### v2

# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

Abstract
Quantifying the complexity of quantum states that possess intrinsic structure, such as symmetry or encoding, in a fair manner constitutes a core challenge in the benchmarking of quantum technologies. This paper introduces the Reference-Contingent Complexity (RCC), an information-theoretic measure calibrated by the available quantum operations. The core idea is to leverage the quantum relative entropy to quantify the deviation of a quantum state from its "structured vacuum"—namely, the maximum entropy state within its constrained subspace—thereby only pricing the process of creating non-trivial information. Our central result is a key theorem that rigorously proves the RCC serves as a lower bound for the complexity of any universal quantum circuit. This lower bound is comprised of a linear dominant term, a universal logarithmic correction, and a precise physical correction term that accounts for non-uniformity in the spectral distribution. Crucially, we establish a set of operational protocols, grounded in tasks like quantum hypothesis testing, which make this theoretical lower bound experimentally "auditable." This work provides a "ruler" for quantum technology that is structure-fair and enables cross-platform comparison, thereby establishing a strictly verifiable constraint between the computational cost of the process and the structured information of the final state.
Keywords Quantum circuit complexity , Structural fairness , Quantum benchmarking , Single-shot quantum information theory , Relative entropy , Kolmogorov complexity
1 Introduction
What, precisely, might constitute the intrinsic complexity of a many-body quantum state? Driven by quantum information science, this fundamental question is increasingly becoming a bridge connecting computational theory and fundamental physics, profoundly influencing our understanding of the potential of quantum computation, the black hole information paradox [1, 2] and its later developments [3], the emergence of spacetime geometry [4, 5], and even phases of matter. At the core of answering this question lies the quantification of the minimum computational resources required to prepare a specific quantum state. This perspective is not only key to assessing the potential of quantum algorithms but also provides a fundamental benchmark for measuring the true performance of quantum hardware, with concrete diagnostics developed in recent works [6, 7]. Although various profound paradigms for complexity measures have been developed in the field—including Nielsen’s complexity based on geometric shortest paths[8], Krylov complexity based on operator algebra growth [9], and resource theories closely related to quantum computational advantage (e.g., magic-state and related approaches [10, 11], recent links to operator growth [12], and general resource-theoretic formalisms [13])—a long-standing challenge remains to be systematically addressed: how to conduct a fair complexity accounting for a “structured” quantum system. The so-called “structural fairness problem” originates from the ubiquitous presence of encodings and symmetries in quantum information processing, which can dramatically reshape what should count as a genuine computational cost [8]. Taking a fault-tolerant quantum computer as an example, its information is encoded within a certain “code subspace” of a vast physical Hilbert space; for a basic operation executed on a logical qubit, its true computational cost should depend on the change in the logical information, rather than the apparent consumption of mobilizing many physical qubits to maintain the error-correcting code structure [14, 15]. In practice, logical operations may be realized via nontrivial transformations supported by the code architecture itself [16, 17]. If a complexity measure cannot distinguish between these two, it could lead to the misleading conclusion that the larger the physical scale of the system and the more complex the encoding structure, the higher its complexity is assessed to be [13]. Similarly, in symmetry-protected topological (SPT) phases or strongly correlated systems, the low-energy physics is confined to specific symmetry sectors; a fair measure ought to treat these structures—endowed “for free” by the Hamiltonian’s symmetries—as an a priori background, and only measure the cost of generating non-trivial quantum correlations on top of this background [18, 19]. This viewpoint is consistent with broader perspectives on emergent phases and constrained dynamics [20]. Existing frameworks for measures often find it difficult to systematically isolate these “free structures,” leading to a lack of a unified and fair benchmark when making cross-comparisons of complexity across different physical systems or encoding schemes [10, 11]. To directly address this challenge, this paper establishes a theory for quantum complexity lower bounds that is structurally fair, theoretically rigorous, and operationally accessible. Its core is the introduction and systematic construction of the theoretical framework of “Reference-Contingent Complexity” (RCC), and its core contributions can be summarized into the following three points:
  1. 1.
    Establishing a structurally fair complexity benchmark: We start from first principles and introduce the mathematical construct of a “reference set,” to formalize any given physical structure, whether originating from symmetries or encoding schemes, as an a priori background. We then define complexity as the information-theoretic distance of a quantum state $\rho$ relative to this background (represented by the maximum-entropy reference state $\sigma_{R}$), namely, the quantum relative entropy $D(\rho||\sigma_{R})$. This construction fundamentally ensures that our proposed measure (RCC) only prices the non-trivial information that transcends the a priori structure, thereby addressing the “structural fairness problem.”
  2. 2.
    Proving it as a theorem-level lower bound for universal circuit complexity: We prove a central theorem establishing that RCC is a rigorous lower bound on the complexity of any universal quantum circuit. This lower bound is dominated by a linear leading term, supplemented by a universal logarithmic correction, and includes a physical correction term $\Delta_{\mathrm{spec}}$ with a coefficient of exactly $1$, which characterizes the spectral distribution of the quantum state. This result establishes a quantitative link between the physical cost of the computational process and the information-theoretic properties of the final state.
  3. 3.
    Connecting the theory with operational observables:To connect this theoretical lower bound with physical reality, we construct a set of operational protocols based on schemes such as quantum hypothesis testing. These protocols transform an abstract theoretical concept into a set of physical observables that can be estimated experimentally with statistical confidence, providing a practical theoretical tool for complexity benchmarking in real quantum systems.
The structure of this paper is organized as follows. We first systematically establish the theoretical framework for Reference-Contingent Complexity (RCC) in Sec. 2, and then proceed in Sec. 3 to propose and prove the central theorem of its role as a strict lower bound for circuit complexity. To connect this theory with physical reality, Sec. 4 delves into the operational principles of RCC. Building on this foundation, Sec. 5 presents the axiomatic properties of RCC and its three major application interfaces, while Sec. 6 discusses it within a broader theoretical context. Sec. 7 provides a summary of the entire paper. To ensure the self-consistency of the manuscript and the completeness of the arguments, all technical derivations are systematically organized in the appendices. Among them, the core theoretical cornerstones—the rigorous proof of the main theorem, the robustness of the framework, and the operational protocols—are elaborated in detail in Appendices A, B, C, and D, respectively. Building upon this, we further extend the boundaries of the theory: Appendices E and F explore the connections of this framework with other theories and its generalization to broader settings, while Appendix G provides constructive evidence for the central conjecture in the main text. Finally, to support the key application interfaces in the main text, Appendix I provides a complete technical supplement on its connection to thermodynamics. A quick reference list of core symbols and terminology can be found in Appendix H.
2 Theoretical Framework: Reference-Contingent Complexity (RCC)
To establish a structurally fair theory of complexity, we must first precisely define the mathematical objects and physical setting that constitute its foundation. This chapter aims to systematically elaborate the core construction of Reference-Contingent Complexity (RCC). We begin from the setting of a universal quantum system, introduce the concept of the “reference set” as the cornerstone of the theory—a mathematical tool to formalize the system’s intrinsic structure as an a priori background—and ultimately, from first principles, provide a complexity measure rooted in information theory and possessing a clear physical meaning.
2.1 System Setting and Notational Conventions
We consider a finite-dimensional quantum system defined on a Hilbert space $H$ of total dimension $d$. The dynamical evolution of the system is driven by a universal finite gate set $G\subseteq\mathrm{SU}(d)$ with cardinality $g$. Our fundamental objective is to quantify the minimum resource cost required to prepare a target state $\rho$ from a simple initial state (such as a product state) using a quantum circuit composed of this gate set. This minimum resource cost, the optimal circuit complexity $C^{*}_{\mathrm{opt}}(\epsilon)(\rho)$, is defined as the length of the shortest sequence of atomic instructions capable of preparing $\rho$ within a precision $\epsilon$ (as measured by the trace distance). To ensure that our complexity benchmark can be compared with concrete physical implementations, we must objectively and unambiguously calibrate the fundamental unit of computational operations. This leads to the core postulate of our framework:
Postulate A (Atomic Information Cost). The fundamental computational cost of any quantum circuit is quantified by the total number of information bits required to describe the shortest corresponding serial sequence of atomic instructions, $C_{\mathrm{opt}}\cdot\log(\Gamma_{R})$, where:
  1. 1.
    $C_{\mathrm{opt}}$ is the length of the shortest sequence of atomic instructions.
  2. 2.
    $\log(\Gamma_{R})$ is the information cost of a single atomic instruction, i.e., the “information bandwidth.” This quantity is uniquely determined by the hardware’s capability, namely the size of the alphabet of atomic instructions, $\Gamma_{R}$.
The value of $\Gamma_{R}$ is based on a rigorously formalized computational model that combines the gate set $g$ with the hardware’s addressability $|S_{R}|$. A complete exposition of this model, including the precise definition of $\Gamma_{R}=g\cdot|S_{R}|$ and the handling of parallel circuits and continuous-parameter gates, is axiomatically defined in detail in Appendix A. This postulate directly links abstract computational steps with the physical capabilities of the hardware (gate set size and addressing range), enabling our theory to capture the essential differences in computational power among different physical platforms. To ensure the self-consistency of the theory, we list here the key information-theoretic quantities that will be used in the subsequent derivations. The von Neumann entropy of a quantum state $\rho$ is defined as $S(\rho)=-\mathrm{Tr}(\rho\cdot\log(\rho))$[21]. The quantum relative entropy between two quantum states $\rho$ and $\sigma$ is $D(\rho||\sigma)=\mathrm{Tr}[\rho\cdot(\log(\rho)-\log(\sigma))]$;[22] on the other hand, throughout this paper, to maintain consistency with information-theoretic conventions, all logarithms are assumed to be base-2 (units in bits) by default. In Sec. 5.4 and Appendix I, where we interface with the framework of thermodynamics, we will primarily use the natural logarithm (units in nats) for convenience of calculation, and will make explicit conversions when necessary.
(Note: To ensure the unambiguousness of the above definitions, this paper adopts the following standard conventions. (i) The parameter $\Gamma_{R}$ only accounts for the size of the discrete gate set alphabet; the precision cost of continuous parameters (e.g., angles of rotation gates) is incorporated into the logarithmic correction term in Theorem (3.1). (ii) The circuit complexity $C_{opt}$ refers to the total number of serial atomic instructions; the parallelism of a physical circuit has been converted into an equivalent serial length. (iii) The core theory of this paper is built upon the unitary circuit model; for the scope of computational models involving measurements and feedback, see Appendix G.3.)
2.2 Reference Set and Structured Priors
The realization of “structural fairness” depends on the precise mathematical characterization of the system’s intrinsic “structure.” The true cost of a logical operation on a fault-tolerant quantum computer should be independent of the vast number of physical qubits mobilized to maintain the error-correcting code; for a system with a conservation law, its dynamical evolution is restricted to specific symmetry sectors, and this constraint is “for free.” To systematically isolate these structures endowed by physical laws or deliberate encoding, we introduce the core construct of this paper.
Definition 2.1 (Constrained Reference Set). A reference set $R$ is a collection of mutually commuting constraint projectors $\{P_{i}\}$, whose common image space is defined by the total projector $\Pi_{R}=\prod_{i}P_{i}$, with the requirement that $\Pi_{R}\neq 0$. This total projector defines the structured subspace $H_{R}=\Pi_{R}\cdot H$, with dimension $d_{R}=\mathrm{Tr}(\Pi_{R})$.
(Note: For structures defined by mutually exclusive sectors (such as eigenspaces of different conserved charges), one should first select the total projector for a target sector, $P_{\mathrm{sector}}$, as $\Pi_{R}$ (or as a member of $\{P_{i}\}$), and then define finer internal structures, so as to avoid the product of multiple mutually exclusive projectors being zero.)
The reference set $R$ physically formalizes the “structured priors” of the system. In a system with a $U(1)$ conserved charge $Q$, one should first fix the total projector for the target sector $P_{\mathrm{sector}}$ (corresponding to a given charge value), and then select mutually commuting internal constraints $\{P_{i}\}$ within that sector, such that $\Pi_{R}=P_{\mathrm{sector}}\cdot(\prod_{i}P_{i})\neq 0$; in a stabilizer error-correcting code, $\{P_{i}\}$ would be the generators of the stabilizer group, and their common +1+1 eigenspace is the code subspace. Once the reference set is given, we can define a state that is most disordered under this structural constraint.
Definition 2.2 (Reference State). The reference state $\sigma_{R}$ associated with a reference set $R$ is defined as the maximally mixed state on the subspace $H_{R}$ defined by the total projector $\Pi_{R}$. Its mathematical form is:
$\sigma_{R}=\frac{\Pi_{R}}{d_{R}}$ (2.1)
The reference state $\sigma_{R}$ physically plays the role of a “zero point in the structured background.” It contains all the structural information encoded by the reference set $R$, but beyond that, it contains no additional quantum correlations or information that would require non-trivial dynamical evolution to generate. Therefore, $\sigma_{R}$ can be regarded as a physically meaningful “structured vacuum” state. The information-theoretic “distance” of a quantum state relative to this vacuum naturally quantifies the non-trivial information content it embodies, which transcends the a priori structure. A detailed discussion on how to objectively select a reference set based on physical principles, and the intrinsic robustness of our framework with respect to this choice, is provided in Appendix F.3.
2.3 Definition of Reference-Contingent Complexity
Based on the foregoing preparations, we can now give the formal definition of Reference-Contingent Complexity. RCC aims to quantify the degree of information “excitation” of a quantum state $\rho$ (whose support lies within the subspace $H_{R}$) relative to its structured background $\sigma_{R}$, normalized by the maximum amount of information that a single universal operation can provide.
Definition 2.3 (Reference-Contingent Complexity): For a quantum state $\rho$ whose support is contained within the subspace $H_{R}$, its Reference-Contingent Complexity $C_{R}(\rho)$ with respect to the reference set $R$ is defined as:
$C_{R}(\rho)=\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$ (2.2)
where $\Gamma_{R}$ is defined by Postulate A. This definition directly links complexity to the distinguishability of two quantum states. A highly complex state must be extremely easy to distinguish from its structured vacuum $\sigma_{R}$. By expanding the relative entropy $D(\rho||\sigma_{R})$ and utilizing the property of the reference state $\sigma_{R}$ as a maximally mixed state, we can obtain an equivalent expression with more physical intuition [^1] It should be emphasized that the strict validity of this equivalence and the finiteness of the relative entropy $D(\rho\|\sigma_{R})$ are conditional on the support of the quantum state being entirely contained within the subspace $H_{R}$, i.e., $\mathrm{supp}(\rho)\subseteq H_{R}$. Considering that unavoidable state leakage in physical systems may cause this condition to be violated, we have discussed in detail in Appendix B.1 how to handle this robustly using mathematical techniques such as projection or smoothing.:
$C_{R}(\rho)=\frac{1}{\log\Gamma_{R}}(\log d_{R}-S(\rho))$ (2.3)
Eq. (2.3) reveals the physical meaning of RCC. Its numerator term, $(\log d_{R}-S(\rho))$, represents the “information deficit” or “degree of order” of the quantum state $\rho$ within its respective structured subspace $H_{R}$. $\log d_{R}$ is the maximum Shannon information capacity of this subspace (i.e., $S(\sigma_{R})$), while $S(\rho)$ is the von Neumann entropy actually contained in the state $\rho$. The difference between the two is precisely the degree to which the state deviates from “complete disorder,” and it is also precisely the amount of non-trivial information that must be “injected” by any quantum circuit that prepares the state. At the same time, the denominator term $\log\Gamma_{R}$ provides a precise calibration of the fundamental unit of computational resources. Therefore, the physical picture of RCC is: the structurally fair complexity of a quantum state is equal to the minimum amount of information required to be injected for its preparation, divided by the maximum information bandwidth that a single complete computational operation can provide. To specifically illustrate how RCC achieves structural fairness, we once again consider a stabilizer code that encodes $k$ logical qubits into $N$ physical qubits. The dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For a logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assuming logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and that a single operation has an addressable range covering all $k$ logical qubits, i.e., $|S_{R}|=k$. According to Eq. (2.3) and the definition $\Gamma_{R}=g_{L}\cdot|S_{R}|$, its RCC is $C_{R}(\rho_{L})=k/\log(g_{L}\cdot k)$. This result aligns with physical intuition: the complexity of this logical pure state is proportional to the number of logical information bits it carries, $k$, and is independent of the number of physical qubits used for encoding, $N$, while the denominator correctly reflects the cost of the logical operation set. This clearly demonstrates how RCC, by its very definition, automatically strips away the ‘free’ encoding structure and only prices the non-trivial logical information. In Appendix C, we provide more detailed implementation guidance by treating this example alongside another core physical scenario (fixed particle number sector) as canonical case studies.It is noteworthy that, from this global perspective, all pure states ($S(\rho)=0$) within the subspace $H_{R}$ share the same maximal leading term $\log(d_{R})/\log(\Gamma_{R})$. This degeneracy is a direct consequence of the information-theoretic nature of RCC. We will revisit this profound feature in Sec. 6.3, where we demonstrate how this degeneracy is lifted by introducing a physically motivated “observational window,” thereby distinguishing the rich internal structures of different pure states.
2.4 A Note on Units: The Structon
To facilitate the understanding of the physical role of the “information bandwidth” $\log\Gamma_{R}$ in Eq. (2.3) and to provide a unified language for comparing complexity across hardware platforms, this paper introduces an auxiliary unit of measurement specific to RCC. The form of Eq. (2.3), $C_{R}(\rho)=D(\rho||\sigma_{R})/\log\Gamma_{R}$, essentially defines a complexity benchmarked against the alphabet size $\Gamma_{R}$ of a single atomic instruction. To make this physical meaning explicit and to simplify reporting, we define the structon, denoted as $\mathrm{st}_{R}$, as the information-theoretic value corresponding to
[…]
It encapsulates the number of discrete options available for a serial atomic instruction on a given platform (i.e., the “alphabet” size $\Gamma_{R}$) into a concise unit. Accordingly, the Reference-Contingent Complexity can be unambiguously reported in structons, with an expression that is independent of the logarithmic base, ensuring the universality of $\mathrm{st}_{R}$:
$C_{R}(\rho)\ [\mathrm{st}_{R}]=\frac{D_{\text{bits}}(\rho||\sigma_{R})}{\log_{2}\Gamma_{R}}=\frac{D_{\text{nats}}(\rho||\sigma_{R})}{\ln\Gamma_{R}}$ (2.5)
The physical interpretation of this unit is: against the background of a given “structured vacuum” $\sigma_{R}$, 1 $\mathrm{st}_{R}$ represents the complexity required to generate a non-trivial structural unit equivalent to the information of a $\Gamma_{R}$-alphabet”. As an independent and self-consistent notational convention, the introduction of $\mathrm{st}_{R}$ does not alter the form or proof of any theorem in this paper; all correction terms in the main theorem, Theorem 3.1 (spectral skew and logarithmic overhead), when measured in bits or nats, are naturally converted to the $\mathrm{st}_{R}$ unit via Eq. (2.5). To maintain compatibility with traditional information-theoretic expressions, the main text will primarily use the dimensionless form $D/\log\Gamma_{R}$; readers may directly interpret this as a physical quantity in units of $\mathrm{st}_{R}$.A minimal example: Consider a platform with $\Gamma_{R}=12$. If an experimental evaluation yields a relative entropy $D(\rho\|\sigma_{R})=\ln(144)$, the leading term of the complexity lower bound is calculated as $CR(\rho)=D_{\mathrm{nats}}/\ln(\Gamma_{R})=\ln(144)/\ln(12)=2~\mathrm{st}_{R}$. Ignoring logarithmic and spectral corrections, one can report a lower bound on the optimal circuit complexity as $C_{\mathrm{opt}}(\rho)\gtrsim 2~\mathrm{st}_{R}$. A standardized report for the leading term itself would be: $CR(\rho)=2~\mathrm{st}_{R}\;[\Gamma_{R}=12]$. To ensure clarity and reproducibility, we recommend following a minimal set of guidelines when reporting results using $\mathrm{st}_{R}$:
  1. 1.
    Reference Selection: Clearly specify the composition of the reference set $R$ and the alphabet parameter $\Gamma_{R}$ (e.g., $\Gamma_{R}=12$, derived from $g=3$ classes of gates acting on $|S_{R}|=4$ addressable locations, counted by ‘serial atomic instructions”), and clarify that the resolution of continuous-parameter gates only enters the logarithmic overhead term and is not counted in $\Gamma_{R}$, and that parallelism is accounted for by equivalent serial length.
  2. 2.
[…]
  3. 3.
    Result Presentation: Present the numerical value of $C_{R}$ (or its lower bound) with the unit $\mathrm{st}_{R}$.
  4. 4.
[…]
(Note: This section, Sec. 2.4, only provides the reporting and conversion rules for $\mathrm{st}_{R}$; its internal logarithmic base convention does not supersede the conventions in other parts of this paper. The logarithmic base throughout this paper defaults to the established convention of each respective section and appendix (such as Sec. 2.1 and Appendix A); this section has explicitly specified $\log_{2}$ and $\ln$ in Eq. (2.4) and Eq. (2.5) to ensure no ambiguity. As a reporting unit, $\mathrm{st}_{R}$ does not change the form of any theorem or proof.)
3 Main Results: A Structurally Fair Lower Bound for Circuit Complexity
Having precisely constructed the theoretical framework for Reference-Contingent Complexity, we now present and prove the central theorem of this paper. This theorem, under a universal model of quantum computation, links a state complexity in an information-theoretic sense (RCC) with a process complexity in a computational science sense (the optimal number of atomic instruction steps) through a rigorous inequality. This not only endows RCC with a solid meaning in computational complexity theory but also provides a structurally fair lower bound for the resource consumption of quantum circuits, derived from fundamental physical principles.
3.1 Main Theorem
We first state the central theoretical result of this paper. The theorem, under ideal conditions, assumes that the support of the target state $\rho$ is strictly contained within the subspace $H_{R}$, i.e., $\mathrm{supp}(\rho)\subseteq H_{R}$. The rigorous proof of this theorem can be found in Appendix A. For the infinitesimal leakage that is unavoidable in physical reality, the theoretical framework remains robust. By applying projection and normalization to the state or smoothing to the reference state, the main theorem can be shown to still hold, with the addition of only controllable logarithmic correction terms, as detailed in Appendix B.1.
Theorem 3.1 (RCC Lower Bound on Circuit Complexity). For a quantum computation model driven by a universal gate set $G$, any quantum circuit that can prepare a quantum state $\rho$ (whose support is located within the subspace $H_{R}$ defined by the reference set $R$) within a precision $\varepsilon$, its minimum number of atomic instruction steps $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ satisfies the following inequality:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})-\frac{c_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]+\frac{1}{\log\Gamma_{R}}\Delta_{\mathrm{spec}}(\rho)$ (3.1)
where $c_{1}$ is a system-independent universal constant of $O(1)$, and the spectral skew term $\Delta_{\mathrm{spec}}(\rho)$ is defined as:
$\Delta_{\mathrm{spec}}(\rho)=S(\rho)-H_{\mathrm{min}}(\rho)$ (3.2)
where $S(\rho)$ is the von Neumann entropy, and $H_{\mathrm{min}}(\rho)=-\log\lVert\rho\rVert_{\infty}$ is the min-entropy, determined by the largest eigenvalue of $\rho$.Eq. (3.1) constitutes the theoretical core of this paper. The right-hand side of the inequality consists of three terms with clear physical roles:
  1. 1.
    Linear Leading Term: $\frac{1}{\log\Gamma_{R}}D(\rho||\sigma_{R})$, which is $C_{R}(\rho)$. It profoundly reveals that the minimum number of computational steps required to prepare a quantum state is, fundamentally, determined by the structurally fair complexity of that state. The more non-trivial information a state contains within its structured vacuum $\sigma_{R}$ (the reference state), the higher the computational cost, in a linear relationship, that any algorithm preparing it must incur.
  2. 2.
    Logarithmic Correction Term: $\frac{c_{1}}{\log\Gamma_{R}}[\dots]$. This term originates from the unavoidable information-theoretic overhead of describing the circuit itself (its length $C_{\mathrm{opt}}$) and the approximation precision ($\varepsilon$). In a prefix-free universal computation framework, the description of a process or an object requires the consumption of logarithmic information resources related to its own complexity or precision requirements. The growth of this term is much slower than that of the leading term and is secondary on a macroscopic scale, but its presence ensures the theoretical rigor of this lower bound.
  3. 3.
    Spectral Skew Correction Term: $\frac{1}{\log\Gamma_{R}}\Delta_{\mathrm{spec}}(\rho)$. This is a key physical correction, with a coefficient of exactly $1$; this is an additive correction with a coefficient of 1 rigorously derived from an entropy identity (Proposition A.4) in Appendix A, rather than an order-of-magnitude estimate. $\Delta_{\mathrm{spec}}(\rho)$ measures the non-uniformity of the quantum state’s spectral distribution. This term enters the inequality with a positive sign, implying that states with more non-uniform spectral distributions—for instance, a mixed state with a few large eigenvalues and a "sea" of small eigenvalues—conversely have a higher lower bound on their preparation complexity. This conclusion aligns with physical intuition: a state with highly concentrated information or a highly ordered structure ($H_{\mathrm{min}}(\rho)$ is much smaller than $S(\rho)$), compared to a state with a more uniform spectral distribution, has a more special structure, and thus requires a more fine-tuned and complex dynamical evolution to be generated. In particular, for any pure state, we have $S(\rho)=H_{\mathrm{min}}(\rho)=0$, hence the spectral skew term $\Delta_{\mathrm{spec}}(\rho)$ vanishes, making the lower bound expression for pure states more compact.
(Note: It is worth noting that Eq. (3.1) is a bootstrap inequality. Its non-recursive, explicit solution (a closed-form lower bound), which is convenient for direct use in data analysis, is derived in detail along with an implementation scheme in Appendix C.)
3.2 Proof Blueprint
The complete and rigorous mathematical proof of Theorem 3.1 is placed in Appendix A. Here, we only sketch the core logic of its proof. This proof chain systematically connects profound results from computational science, algorithmic information theory, and quantum information theory, and is divided into four steps:
  1. 1.
    Encoding Argument ($C_{\mathrm{opt}}\to K_{\varepsilon}(\rho|R)$): Any quantum circuit composed of $C_{\mathrm{opt}}$ atomic instruction steps can itself be viewed as a computer program that describes $\rho$. Within a prefix-free description framework, the total number of information bits required to describe this circuit (upper-bounded by $C_{\mathrm{opt}}\cdot\log\Gamma_{R}$ plus logarithmic corrections) necessarily constitutes an upper bound on the smoothed relative Kolmogorov complexity $K_{\varepsilon}(\rho|R)$ of the generated state $\rho$.
  2. 2.
    Algorithmic Information Theory Bound ($K_{\varepsilon}(\rho|R)\to D_{\max}^{\varepsilon}$): A fundamental result in quantum algorithmic information theory is that the smoothed Kolmogorov complexity of a quantum state, $K_{\varepsilon}(\rho|R)$, is lower-bounded by its smoothed max-relative entropy $D_{\max}^{\varepsilon}(\rho||\sigma_{R})$ with respect to the reference state $\sigma_{R}$ [23, 24, 25, 26, 27]. This step connects an uncomputable quantity, $K_{\varepsilon}$, with a rigorously defined information-theoretic quantity, $D_{\max}^{\varepsilon}$. (Note: In particular, a key technical point is that the universal constant in this bound is independent of the system’s dimension, thereby ensuring the universality of our final lower bound. Its rigorous proof can be found in Appendix B.2.)
  3. 3.
    Proposition 3.1 (Entropy Identity): For a quantum state $\rho$ with support contained in $H_{R}$, the following identity holds: $D_{\max}(\rho||\sigma_{R})-D(\rho||\sigma_{R})=S(\rho)-H_{\mathrm{min}}(\rho)=\Delta_{\mathrm{spec}}(\rho)$ This identity precisely connects the language of algorithmic information theory with that of standard quantum information theory, with the difference being exactly equal to the spectral skew term.
  4. 4.
    Combining and Rearranging: By chaining together the inequalities and the identity from the three steps above, we obtain a complete logical chain from $C_{\mathrm{opt}}\cdot\log\Gamma_{R}$ to $D(\rho||\sigma_{R})$. After algebraic rearrangement and separation of the leading and correction terms, the final form of Theorem 3.1 is obtained.
3.3 Smoothed Theorem and Operational Connection
Any computational process in physical reality is accompanied by a finite precision $\varepsilon$, and any statistical test operates under a finite confidence level $\eta$. To precisely capture this reality in theory and to establish a more direct connection with operational quantities, we introduce a “smoothed” version of the main theorem. This version adopts the language originating from one-shot quantum information theory, replacing standard information-theoretic quantities with entropy measures that are more physically robust.
Theorem 3.2 (Smoothed Form of the RCC Lower Bound). Under the same conditions as in Theorem 3.1, for any confidence level $\eta\in(0,1)$ and approximation precision $\varepsilon$, there exist system-independent universal constants $c^{\prime}_{1},c^{\prime}_{2}=O(1)$ such that:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D_{H}^{\eta}(\rho||\sigma_{R})-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\log(1/\eta)$ (3.3)
The core significance of Theorem 3.2 lies in the fact that it connects the lower bound of complexity to a physical task with the most direct operational definition: distinguishing the quantum state $\rho$ from its structured vacuum (the reference state) $\sigma_{R}$ with an optimal strategy. The fundamental difficulty of this task is quantified by the hypothesis testing relative entropy, $D_{H}^{\eta}(\rho||\sigma_{R})$. This quantity precisely characterizes the minimum achievable exponential decay rate of the type-II error probability, under the condition that the type-I error (mistaking $\sigma_{R}$ for $\rho$) probability does not exceed $\eta$. To rigorously establish this connection, we provide a complete proof in Appendix A.6. This proof systematically employs core tools from one-shot information theory, such as the smoothed max-relative entropy, and ultimately unifies the physical cost of a computational process, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, with the statistical limit of an information-theoretic task, $D_{H}^{\eta}(\rho||\sigma_{R})$, through a rigorous chain of inequalities.[21] This ultimately proclaims that RCC is not only a self-consistent theoretical construct but also a quantity that can, in principle, be probed through physical measurements. This theorem not only enhances the robustness of our theoretical framework, but more importantly, it paves a solid theoretical path for the operational schemes proposed in Chapter 4 for estimating the complexity lower bound based on statistical experiments. The standard information-theoretic relations connecting the entropy measures used in this theorem and in Theorem 3.1 are provided in Appendix C.
4 Operational Principles and Information Extraction for RCC
The vitality of any profound physical quantity lies not only in its theoretical self-consistency and elegance, but more so in its connection to observables. The preceding sections have established the theoretical framework of Reference-Contingent Complexity and have proven its central role as a lower bound on the optimal gate count. However, if RCC itself could not be probed by any physical means, this lower bound would lose its practical significance. This chapter aims to bridge this gap between theory and reality by systematically elaborating on several fundamental principles for extracting RCC information from a physical system, thereby transforming an abstract theoretical concept into a set of concrete physical measurement protocols with statistical confidence. This chapter will elaborate on these fundamental principles separately, while a practical protocol that integrates them into a unified operational workflow is provided in Appendix B.3.2. On the other hand, this chapter will primarily discuss protocols in ideal finite-dimensional systems, whereas a rigorous framework for robust handling and conservative lower-bound schemes for the unavoidable energy or mode cutoffs when dealing with infinite-dimensional systems (such as quantum field theories) is detailed in Appendix F.
4.1 Principle A: Quantum Hypothesis Testing
From the fundamental perspective of information theory, the distinguishability between two quantum states is characterized by the theory of hypothesis testing. Theorem 3.2 has already revealed that the smoothed lower bound of RCC is directly related to the hypothesis testing relative entropy $D_{H}^{\eta}(\rho||\sigma_{R})$. This provides us with the first, and theoretically most fundamental, principle of information extraction. Its core task can be formulated as a binary quantum hypothesis testing problem: given a single-copy system, determine whether its state satisfies the null hypothesis $H_{0}:\sigma_{R}$ or the alternative hypothesis $H_{1}:\rho$. Any given measurement (described by a POVM $\{T,I-T\}$) will yield a lower bound on $D_{H}^{\eta}(\rho||\sigma_{R})$, which alleviates the stringent requirement of finding the optimal measurement, making this principle feasible in practice. At the implementation level, let us assume $N$ independent and repeated measurements are performed, from which the frequency of the type-II error (mistaking $\rho$ for $\sigma_{R}$) is observed, and its probability $\beta$ is thereby estimated. To ensure statistical rigor, we need to construct a confidence interval for $\beta$. Using standard statistical methods, such as the Hoeffding bound or the more precise Clopper-Pearson interval [28, 29], one can calculate the upper confidence endpoint of $\beta$, denoted $\beta_{U}(\delta)$, at a confidence level of $1-\delta$. (Note: In the chosen binary test $\{T,I-T\}$, it is specified that observing $T$ leads to the decision $H_{1}(\rho)$, while observing ITI-T leads to the decision $H_{0}(\sigma_{R})$. The type-I error is constrained as $\alpha=\mathrm{Tr}[T\cdot\sigma_{R}]\leq\eta$. With a finite sample size, the upper confidence endpoint of $\alpha$, $\alpha_{U}(\delta)$, can be obtained using the Clopper–Pearson (or Wilson) interval, and it should be ensured that $\alpha_{U}(\delta)\leq\eta$; under this constraint, the lower bound of $D_{H}^{\eta}(\rho||\sigma_{R})$ is then assessed based on the upper confidence endpoint of $\beta$, $\beta_{U}(\delta)$.) We then have a statistical lower bound for the leading term of RCC:
Proposition 4.1 (Hypothesis Testing Lower Bound):
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log\beta_{U}(\delta)$ (4.1)
This inequality reveals a profound physical picture: the intrinsic complexity of a quantum state $\rho$ is lower-bounded by our ability to confidently distinguish it from its “structured vacuum” $\sigma_{R}$ in an experiment. A highly complex state must have extremely high distinguishability, corresponding to an extremely low type-II error rate, which in turn implies a large lower bound on its complexity. The complete experimental protocol, including the rigorous control of statistical errors, is detailed in Appendix D.1. Of particular importance is that the advantage of the present theoretical framework lies in the fact that this connection between “distinguishability” and “complexity” does not merely remain at the level of physical intuition. It is elevated by Theorem 3.2 to a mathematically provable relationship. It is this theorem, along with its complete derivation based on one-shot information theory in Appendix A.6, that builds the bridge from experimental data (the type-II error rate $\beta$) to the theoretical lower bound ($C_{\mathrm{opt}}$). This rigorous connection ensures that our proposed operational schemes are theoretically completely reliable and self-consistent, successfully transforming an abstract concept of complexity into a measurable physical reality.
4.2 Principle B: Projective Witnesses
The proof chain of the main theorem reveals a profound connection between RCC and the max-relative entropy, $D_{\max}$. We can probe this relationship by measuring a carefully chosen projection operator, thereby providing a rigorous lower bound for $D_{\max}$, and in turn, for the complexity itself.
Lemma 4.1 (Projective Occupation Lower Bound): Consider a projection operator $\Pi$ of rank $r$, whose support lies within the subspace $H_{R}$. Its occupation probability on the state $\rho$ is $p=\mathrm{Tr}(\Pi\cdot\rho)$. Then $D_{\max}(\rho||\sigma_{R})$ satisfies:
$D_{\max}(\rho||\sigma_{R})\geq\log(p\cdot d_{R}/r)$ (4.2)
Proof. If $\rho\leq 2^{\lambda}\cdot\sigma_{R}$, then for a projector $\Pi$ of rank $r$, taking the trace yields $\mathrm{Tr}(\Pi\cdot\rho)\leq 2^{\lambda}\cdot\mathrm{Tr}(\Pi\cdot\sigma_{R})=2^{\lambda}\cdot r/d_{R}$. From $p=\mathrm{Tr}(\Pi\cdot\rho)$, one can solve for $\lambda\geq\log(p\cdot d_{R}/r)$. The lemma is thus proven. In experimental practice, we estimate the value of $p$ through $N$ measurements and calculate its lower confidence endpoint $p_{\mathrm{lower}}(\delta)$ at a confidence level of $1-\delta$. Substituting this statistical lower bound into Eq. (4.2) yields a lower bound for $D_{\max}$ with statistical confidence. The physical meaning of this result is exceptionally clear: if it is found experimentally that the quantum state $\rho$ occupies, with a high probability $p_{\mathrm{lower}}$, a subspace (defined by $\Pi$) of dimension $r$ much smaller than the dimension $d_{R}$ of its containing structured subspace, then this serves as a proof that the information in this state is highly “concentrated” or “ordered.” This high concentration of information implies that the state $\rho$ possesses a very large eigenvalue, and thus $H_{\mathrm{min}}(\rho)$ is small and $D_{\max}$ is large. Therefore, through a simple projective measurement, we have “witnessed” the non-triviality of the state, setting a rigorous lower bound on its complexity. The detailed implementation protocol for this principle, including the statistical estimation of occupation probabilities and confidence level corrections in scenarios with multiple witnesses, is formally articulated in Appendix D.2.
4.3 Principle C: Dephasing Information Bound
The third principle extracts information through a “classicization” process. To obtain a reliable lower bound, we employ the completely dephasing channel $\Delta_{R}$ derived from the reference set $R$, which corresponds to performing a measurement in a basis of rank-1 projectors that simultaneously diagonalizes all projection operators $\{P_{i}\}$ and then discarding the off-diagonal (coherent) information. Physically, this channel corresponds to the strongest decoherence process that leaves the structure $R$ invariant.[21] According to the property of entropy increase in quantum information theory, the von Neumann entropy after the action of this channel is necessarily no less than before, i.e., $S(\Delta_{R}(\rho))\geq S(\rho)$. The dephased state $\Delta_{R}(\rho)$ is a classical probability distribution, and its Shannon entropy is denoted as $H=S(\Delta_{R}(\rho))$. Therefore, we have $H\geq S(\rho)$, which provides an operational and conservative lower bound for RCC:
Proposition 4.2 (Dephasing Lower Bound):
$C_{R}(\rho)\geq\frac{(\log d_{R}-H)}{\log\Gamma_{R}}$ (4.3)
Experimentally, this corresponds to performing measurements in a classical basis compatible with the reference set $R$, reconstructing the probability distribution, and estimating its Shannon entropy $H$. An upper confidence interval for $H$ can be given through the DKW inequality or bootstrap methods,[30] thereby propagating the statistical uncertainty to the final lower bound of RCC. The detailed implementation protocol for this principle, particularly concerning how to rigorously estimate Shannon entropy from finite samples with bias correction, is presented with its complete statistical framework in Appendix D.3. Furthermore, the three principles proposed in this chapter can be integrated into an operational pipeline with statistical confidence, as detailed in Appendix B.
4.4 An Important Clarification: The Role of Purity
In complexity measures, the role played by purity, $P=\mathrm{Tr}(\rho^{2})$, is often misunderstood. A measurement of purity provides an upper bound on $S(\rho)$, not a lower bound, because the von Neumann entropy $S(\rho)$ is always greater than or equal to the second-order Rényi entropy $S_{2}(\rho)=-\log P$. Therefore, substituting this relationship into the expression for RCC, Eq. (2.3), a purity measurement can only provide an upper bound on complexity: $C_{R}(\rho)\leq\frac{(\log d_{R}+\log P)}{\log\Gamma_{R}}$. It cannot be used to prove a lower bound on complexity; its more appropriate role is as a sanity check tool. A correct understanding is that “high purity” in itself does not guarantee a large RCC lower bound. A meaningful lower bound must originate from the distinguishability of the state $\rho$ relative to the reference state $\sigma_{R}$. This distinguishability is quantified by the success rate of hypothesis testing, a high occupation probability on a low-rank projection, or a classical entropy that is far from the uniform distribution, rather than by purity itself. In contrast, the three operational principles proposed in this chapter can all provide rigorous lower bounds. Detailed protocols for the three principles mentioned above can be found in Appendix D, and their technical details and boundary conditions on issues such as multiple testing, statistical power control, and readout leakage propagation should be consulted in conjunction with Appendix C.
5 Properties and Application Interfaces of RCC
The preceding sections have established the definition of Reference-Contingent Complexity (RCC), its theorem as a lower bound for optimal gate count, and the operational principles for extracting its information. A well-behaved physical quantity should not only be operational but also adhere to a series of universal physical principles and be able to seamlessly interface with existing theoretical frameworks. This chapter aims to reveal that RCC indeed possesses such qualities. We will first systematically elaborate on the axiomatic properties that RCC satisfies as a well-behaved resource measure, and subsequently showcase its three key application interfaces: first, we will elucidate how the introduction of RCC deepens and unifies the performance frontier of quantum dynamics; second, how it provides a structurally fair calibration scheme for exploring fundamental questions such as the complexity-geometry duality; and finally, this paper will reveal its third application as a tool for quantifying thermodynamic order and constraining statistical process laws, demonstrating how this state function can provide rigorous lower bounds for the information processing work and evolution time in physical processes.
5.1 Axiomatic Properties as a Resource Measure
A measure for a physical resource should exhibit good and consistent behavior within the theoretical framework it defines. We will show that RCC satisfies a series of core axiomatic properties of a "well-defined" physical quantity, with most of these favorable properties being directly inherited from quantum relative entropy, which serves as its foundation.
Proposition 5.1 (Axiomatic Properties of RCC): For a given reference set $R$ and information bandwidth $\log\Gamma_{R}$, the Reference-Contingent Complexity $C_{R}(\rho)$ satisfies the following properties:
  1. 1.
    Monotonicity under cluster-preserving channels: For a class of "free operations" that preserve the structure of the reference set [^2] Here, ”free operations” are formally defined as a class of cluster-preserving CPTP maps $E$ that maintain the structure of the reference set, with their core property being $E(\sigma_{R})=\sigma_{R}$. The monotonicity of RCC under this class of operations, $C_{R}(E(\rho))\leq C_{R}(\rho)$, is a direct corollary of the data processing inequality for quantum relative entropy, $D(E(\rho)\|E(\sigma_{R}))\leq D(\rho\|\sigma_{R})$.—that is, CPTP maps $E$ satisfying $E(\sigma_{R})=\sigma_{R}$—RCC is necessarily non-increasing:
    $C_{R}(E(\rho))\leq C_{R}(\rho)$ (5.1)
    This property originates from the data processing inequality for quantum relative entropy, $D(E(\rho)||E(\sigma_{R}))\leq D(\rho||\sigma_{R})$, and it ensures the self-consistency of RCC as a resource that cannot be created "for free," fully conforming to the general framework of quantum resource theories [13]. (Note: A rigorous definition of "free operations" and the objectivization criteria that should be followed for selecting a physically appropriate reference set are discussed in depth in Appendix F.3.)
  2. 2.
    Convexity: RCC is a convex function of the state $\rho$. For $\rho=\sum_{i}p_{i}\cdot\rho_{i}$, where $\{p_{i}\}$ is a probability distribution, we have $C_{R}(\rho)\leq\sum_{i}p_{i}\cdot C_{R}(\rho_{i})$. This implies that preparing a new state by incoherently mixing multiple states does not result in a complexity exceeding the weighted average of the complexities of its components, reflecting the reduction in order due to information mixing.
  3. 3.
    Unitary invariance: For any unitary operation $U$ that preserves the subspace $H_{R}$ (i.e., $U\cdot H_{R}\cdot U^{\dagger}=H_{R}$), RCC remains unchanged due to the unitary invariance of the von Neumann entropy: $C_{R}(U\cdot\rho\cdot U^{\dagger})=C_{R}(\rho)$.
  4. 4.
    Extrema and bounds: RCC has a well-defined range of values, $0\leq C_{R}(\rho)\leq\log(d_{R})/\log\Gamma_{R}$. Its lower bound is achieved when $\rho=\sigma_{R}$ (i.e., the reference state itself), and its upper bound is achieved when $\rho$ is any pure state within $H_{R}$.
  5. 5.
    Conditional monotonicity under cluster refinement: If the reference set is refined from $R$ to $R^{\prime}$ ($R\subset R^{\prime}$), leading to a stronger structural constraint ($H_{R^{\prime}}\subset H_{R}$), then:
    1. (a)
      a. If the hardware addressability remains unchanged ($\Gamma_{R^{\prime}}=\Gamma_{R}$), the complexity does not increase: $C_{R^{\prime}}(\rho)\leq C_{R}(\rho)$.
    2. (b)
      b. If the stronger constraint limits the addressability ($\Gamma_{R^{\prime}}<\Gamma_{R}$), then monotonicity is not guaranteed.
In summary, these favorable mathematical properties collectively demonstrate that RCC is a theoretically self-consistent and well-behaved physical quantity, enabling it to serve as a stable benchmark for comparison and analysis in various physical contexts.
5.2 Application Interface 1: Providing an Operational Lower Bound for the Dynamical Performance Equation
As an independent and self-consistent lower bound on complexity, an important application of RCC is to provide an operational interface for our previous work on the ’energy-entanglement performance equation’ [31][32, 33]. In that work, by unifying quantum speed limits and principles of locality, we established an "energy-entanglement performance equation" (RECT-$\eta$) that connects dynamical resources, information output, and the optimal computational complexity $C_{\mathrm{opt}}$. The core conclusion of that framework is a precise ’resource accounting identity’ (its full derivation is presented in Appendix E), which directly relates physical quantities of the system to theoretical complexity:
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (5.2)
where $\eta_{\mathrm{QSL}}$ and $\eta_{\mathrm{LR}}$ are dynamical factors that quantify the efficiency of the system’s utilization of energy resources and entanglement propagation, respectively; $\gamma$ is the locality constant of the LR bound (dimensionless); and $\sigma_{\mathrm{avail}}$ represents the available energy fluctuation / effective energy bandwidth (a process-dependent energy scale). Although Eq. (5.2) is theoretically universal, the $C_{\mathrm{opt}}$ it contains is a theoretical limit that cannot be directly measured, which restricts the direct experimental application of the framework. The RCC framework proposed in this work provides a solution to this. By applying our core result, Theorem 3.1, which provides a rigorous lower bound for $C_{\mathrm{opt}}$ with $C_{R}(\rho)$ (i.e., $C_{\mathrm{opt}}\geq C_{R}(\rho)$), we can transform the above theoretical identity into a performance inequality that can be directly used for experimental and numerical analysis:
$\sigma_{\mathrm{avail}}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{R}(\rho)$ (5.3)
(Note on dimensional analysis: $S_{E}$ is dimensionless, $\sigma_{\mathrm{avail}}$ and $J$ are energy scales, so the units on both sides are consistent; $\hbar$ cancels out during the equation’s derivation.)
To obtain a more universal, dimensionless expression that facilitates cross-platform comparisons, we divide both sides of Eq. (5.3) by the system’s characteristic energy scale $J$, yielding:
$\frac{\sigma_{\mathrm{avail}}}{J}\cdot S_{E}\geq\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma}{2}\right)\cdot C_{R}(\rho)$ (5.4)
Eq. (5.4) represents a significant conceptual deepening. The introduction of RCC transforms the core complexity of the original framework from an abstract quantity $C_{\mathrm{opt}}$, which is dependent on the specific evolution path and is not measurable, into a state function $C_{R}$, which is determined solely by the final state $\rho$ of the system and is, in principle, measurable. This reveals a more profound and more operational physical constraint: regardless of the ingenious path the system evolves through, the final achievable "resource-output product" (the product of energy fluctuation normalized by the characteristic energy scale and the final state’s entanglement entropy) is strictly lower-bounded by the structurally fair complexity $C_{R}$ inherent in the final state itself. Therefore, RCC plays a key bridging role here, transforming a purely theoretical relation (the RECT-$\eta$ identity) into a falsifiable physical proposition directly related to the properties of the final state, thereby greatly expanding the applicability of the original framework. The complete derivation of the aforementioned performance equation, including its profound physical meaning as a "resource accounting identity," is detailed in Appendix E.
5.3 Application Interface 2: Providing a Calibration Scheme for the Complexity-Geometry Duality
In frontier fields such as holographic gravity, conjectures like "complexity = geometry" (e.g., complexity = volume or complexity = action) are highly inspiring, but one of their core obstacles is that the definition of complexity on the boundary field theory side is ambiguous and non-unique [34]. The "structural fairness" idea of RCC provides a "bottom-up," physically well-defined proposal for a structured benchmark. We propose a testable calibration procedure to explore this dual relationship:
  1. 1.
    Determine the structured priors of the boundary theory: On the side of the boundary conformal field theory (CFT), select a family of states with a well-defined symmetry (e.g., $U(1)$ charge conservation). Formalize this symmetry structure as an a priori "reference set" $R$, thereby decomposing the total Hilbert space into independent superselection sectors.
  2. 2.
    Calculate the structurally fair complexity: For this family of states, calculate their structurally fair complexity $C_{R}(\rho)$ relative to the selected sector. When dealing with infinite-dimensional systems such as CFTs, this scheme needs to be complemented with an energy or mode cutoff $\Lambda$ to define the reference state and the subspace dimension $d_{R}(\Lambda)$.
  3. 3.
    Test for an order-preserving relationship with the geometric quantity: On the gravity dual side, calculate the geometric quantities corresponding to these states, such as the wormhole volume $V(\rho)$ or the action of the Wheeler-DeWitt patch $A(\rho)$. Through a normalized mapping, test whether there exists a stable, monotonic, order-preserving relationship between $C_{R}(\rho)$ and the geometric quantity.
In this context, the comparison should primarily focus on order-preservation and scaling behavior, rather than insisting on a direct match of absolute values, to avoid strong dependence of the results on the cutoff method. The advantage of RCC is that it provides a structurally normalized complexity benchmark with a clear physical meaning. If a certain geometric candidate quantity is found to systematically deviate from the order relationship with RCC across different families of states or under different symmetry constraints, it may suggest that the geometric quantity fails to fully capture the physical essence of complexity, or that there is a structural bias in the duality proposal that needs to be corrected. Therefore, RCC is expected to play a key role as a calibration tool in the future construction of a more precise "complexity-geometry" dictionary. (We provide a rigorous framework in Appendix F for calculating a controllable, conservative lower bound on complexity in the presence of energy or mode cutoffs.)
5.4 Application Interface 3: Quantifying Thermodynamic Order and Constraining Process Laws
Following the preceding interfaces that connect RCC with the system’s dynamics (Sec. 5.2) and spacetime geometry (Sec. 5.3), this section aims to reveal its third application: as a tool for quantifying thermodynamic order and constraining statistical process laws. A central challenge in physics is how to use easily measurable "state functions" (such as entropy, free energy) to constrain hard-to-track "process quantities" (such as work, heat, evolution time). This interface aims to provide a novel solution to this challenge. We will demonstrate that two core thermo-dynamical resources consumed by a quantum process, Information Processing Work and Evolution Time, can be rigorously lower-bounded by an operational state function determined solely by the initial and final states of the system, namely, Reference-Contingent Complexity (RCC). All technical derivation details are consolidated in Appendix I. To achieve the above goal, we leverage the theoretical framework of "Complexity Window Thermodynamics" (CWT) [35]. CWT generalizes statistical mechanics to handle information complexity by introducing a finite "complexity budget" $\Xi$. Its core physical quantities include the windowed entropy $S_{\Xi}$, effective temperature $T_{\Xi}$, and information processing work $\Pi_{\Xi}$. To work within this framework, we need to naturally generalize RCC to a quantity that depends on the observational capability $\Xi$. In the following, the partial derivative with respect to $\Xi$, unless otherwise specified, is assumed to be taken at a fixed energy constraint $E$ and is denoted as $(\partial/\partial\Xi)_{E}$.
Definition 5.2 (Windowed Reference-Contingent Complexity). For a given quantum state $\rho$, reference set $R$, and complexity budget $\Xi$, its windowed Reference-Contingent Complexity is defined as
$C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (5.5)
where $S_{\Xi}(\rho)=S(E_{\Xi}(\rho))$ is the windowed entropy of $\rho$ under the conditional expectation map $E_{\Xi}$, and $S$ is the von Neumann entropy. Physically, $C_{R,\Xi}(\rho)$ quantifies the non-trivial degree of order exhibited by the system $\rho$ beyond its a priori structure, under a finite observational capability $\Xi$. As proven in Appendix I.2, $S_{\Xi}$ is monotonically non-increasing with $\Xi$, and therefore $C_{R,\Xi}$ is monotonically non-decreasing with $\Xi$; in the limit $\Xi\to\infty$, it recovers the standard RCC of Sec. 2.3. This generalization is central to connecting the two theoretical frameworks, as it allows us to relate the process quantities in CWT to the state function RCC.
Proposition 5.3 (Work–Complexity Identity). In a physical process that keeps the energy $E$ constant, the complexity generation potential $\Pi_{\Xi}$ (i.e., the information processing work per unit of complexity budget) can be precisely expressed by the rate of change of the windowed RCC:
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (5.6)
Proof Outline. The result is obtained by directly differentiating the definition of $C_{R,\Xi}$ and combining it with the definition of $\Pi_{\Xi}$ in CWT; its rigorous derivation, including the more general case of a variable bandwidth $\Gamma_{R}=\Gamma_{R}(\Xi)$, is detailed in Appendix I.3. This identity serves as a bridge between thermodynamics and information theory: the energy cost to increase the system’s resolvable complexity ($\Pi_{\Xi}$) is proportional to the product of the system’s energy scale ($T_{\Xi}$) and the complexity gain with an increasing budget $($$\partial C_{R,\Xi}/\partial\Xi)_{E}$), with the proportionality constant being the hardware’s information bandwidth ($\log\Gamma_{R}$). Rigorously, the total information processing work is $W_{\mathrm{info}}=\int\Pi_{\Xi}\mathrm{d}\Xi=\log\Gamma_{R}\cdot\int T_{\Xi}\mathrm{d}C_{R,\Xi}$. When $T_{\Xi}$ varies slowly within the interval, $W_{\mathrm{info}}\approx\bar{T}_{\Xi}\cdot\log\Gamma_{R}\cdot\Delta C_{R,\Xi}$, where $\bar{T}_{\Xi}$ is the weighted average temperature along the complexity coordinate (its rigorous definition is given in Appendix I.3.1). This result directly links the macroscopic process quantity $W_{\mathrm{info}}$ with the net change in a state function, $\Delta C_{R,\Xi}$, which is determined solely by the information structure of the initial and final states, providing a generalization of Landauer’s principle [36] for "structured information." In addition to constraining the energy cost, RCC also sets a fundamental lower limit on the time cost of a quantum process.
Theorem 5.4(RCC-Driven Lower Bound on Process Time). Consider a quantum process evolving from an initial state $\rho_{i}$ to a final state $\rho_{f}$ over a total duration $t_{f}$, with an initial optimal complexity $C^{*}(0)$. If the optimal complexity evolution envelope $C^{*}(t)$ of the process adheres to the universal rate upper bound of CWT, $\dot{C}^{*}\leq(2/\hbar)\Pi_{\Xi}$ [35] Prop. 6.1, then
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (5.7)
where $\langle\Pi_{\Xi}\rangle_{t}=(1/t_{f})\int_{0}^{t_{f}}\Pi_{\Xi}(t)\mathrm{d}t$ is the time average of the complexity generation potential. Proof Outline. The result is obtained by integrating the rate bound over time and combining it with the main theorem $C^{*}(t_{f})\geq C_{R}(\rho_{f})$; its rigorous derivation, including full correction terms, is detailed in I.4.
Corollary 5.5 (Lower Bound on Time for Net Process Gain). In scenarios where the net complexity gain of the process is of primary interest,
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (5.8)
This holds strictly when either of the following physical conditions is met:
  1. 1.
    Compact initial state: $C^{*}(0)=C_{R}(\rho_{i})$ (e.g., $\rho_{i}=\sigma_{R}$);
  2. 2.
    Non-retreating envelope assumption: $\Delta C^{*}\geq\Delta C_{R}$.
If the process is approximately isothermal, and one adopts the chaos-bound analogy from CWT, $\Pi_{\Xi}\lesssim\pi k_{B}T_{\Xi}$, then $t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$, which formally resonates with the chaos time scale bound of Maldacena et al. [37]. This result is complementary to the RECT-$\eta$ framework of Sec. 5.2, indicating that the time cost of a quantum process is ultimately constrained by the information-theoretic complexity it must generate. The introduction of RCC allows us to establish a bridge, unified by "information structure complexity," across the three major physical domains of dynamics (Sec. 5.2), geometry (Sec. 5.3), and thermodynamic statistics (this section). The RCC terms in Eqs. (5.7) and (5.8) can, in principle, be experimentally estimated via the operational protocols proposed in Chapter 4; a robustness framework for handling real-world non-idealities (such as state leakage and statistical errors) is provided in Appendix I.5. This paves the way for experimentally testing these fundamental "time-work-information" trade-off relations in real physical systems, and suggests that the computational complexity of a quantum state may not be merely a metric in computational science, but rather a deeper organizing principle of the material world that unifies its various physical manifestations.
6 Discussion
Having systematically constructed the framework of Reference-Contingent Complexity (RCC), proven its theoretical properties, and articulated its application interfaces in the preceding sections, we now place RCC within the broader landscape of theoretical physics for examination. This chapter aims to clarify the relationship between RCC and other relevant complexity measures in the field, frankly discuss the applicability boundaries and open questions of the current theoretical framework, and, based on this, provide an outlook on potential future research directions.
6.1 Comparison and Connection with Related Complexity Measures
RCC is not intended to replace existing complexity measures, but rather to provide a complementary tool with the advantage of structural fairness, which can serve as a universal benchmark. Its relationship with other major paradigms is systematically compared in detail in Appendix B.3; here, we only briefly articulate the core logical relationships.
  1. 1.
    Relationship with Nielsen’s Geometric Complexity: Nielsen complexity [38], from the perspective of differential geometry, defines complexity as the length of the geodesic connecting an initial state to a target state on the manifold of the unitary group, where penalty factors are used to suppress evolution in "difficult" directions. RCC forms a profound complementary relationship with it. Regardless of how the penalty factors are chosen, our main theorem, Theorem 3.1, can provide a rigorous lower-bound anchor point for any geometric path, derived from the fundamental principles of information theory. In other words, the length of any geometric path, after being converted into the number of atomic instruction steps, must be greater than or equal to the lower limit set by the RCC of the final state. This allows RCC to serve as a universal tool for the calibration and sanity-checking of geometric complexity.
  2. 2.
    Relationship with Krylov Complexity: Krylov complexity [39] measures complexity by analyzing the expansion of an operator in the Lanczos/Krylov basis generated by the Hamiltonian and the initial operator, and is closely related to quantum chaos and operator growth. The two focus on different physical aspects: Krylov complexity reflects the operator spreading and chaotic properties under continuous-time evolution, whereas RCC is rooted in the computational framework of discrete gate models and the system’s structural priors. An interesting research direction is to explore the quantitative relationship between RCC and Krylov complexity in the specific setting where the energy eigenbasis is taken as the reference set $R$. We conjecture that for certain chaotic systems, the two may exhibit a monotonic or order-preserving correspondence, which is worthy of systematic numerical and analytical verification in the future.
  3. 3.
    Relationship with Resource Theories such as "Mana": Resource theories represented by "mana" [40] aim to quantify non-stabilizer states, a key resource for achieving universal quantum computation. RCC and this type of specific resource theory form a complementary "cost-benefit" perspective: a high "mana" value is often associated with the ability to perform complex computational tasks, while our work shows that preparing such a high-"mana" state must itself incur a corresponding complexity cost in terms of RCC. RCC provides a unified, information-theory-based "manufacturing cost" pricing for all these specific quantum resources.
6.2 Limitations and Open Questions
The rigor of the RCC framework is built upon several key assumptions, which also define its applicability boundaries and the open questions that remain to be explored. It is worth emphasizing that, within a fixed subspace $H_{R}$, all pure states have the same leading term for RCC, $\log(d_{R})/\log(\Gamma_{R})$. As an information-theoretic lower bound, RCC aims to set a baseline for resource consumption, rather than to distinguish between quantum states that have the same spectral distribution but different geometric or entanglement structures—a common feature of all state functions based on von Neumann entropy. The most central theoretical assumption lies in the strict commutativity of the reference set $R$, which makes it naturally suited for systems defined by Abelian group symmetries or stabilizer subgroups. For systems defined by non-Abelian group symmetries, a possible path for generalization is as follows: technically, one can decompose the Hilbert space into a direct sum of irreducible representations (irreps) according to group representation theory, and use the normalized equally-mixed state of each irrep block (or its energy-truncated version) as the reference state; concurrently, define the CPTP maps that preserve this algebraic structure (such as the conditional expectation onto the center of the algebra) as free operations. (Note: The complete formal construction of this generalization scheme, including the definition of its reference state and the preservation of core properties, has been detailed in Appendix F.1.) Within this framework, the monotonicity of RCC is expected to be preserved based on the data processing inequality, but its proof as a lower bound on circuit complexity would need to be reconstructed. Another core open question is the near-tightness of RCC as a lower bound. In Eq. (3.1), we proved that RCC is a rigorous lower bound for $C_{\mathrm{opt}}$, but this does not answer whether $C_{\mathrm{opt}}$ can also be bounded from above by RCC. If such an upper bound exists, it would imply that RCC is not just a theoretical baseline, but a benchmark with strong predictive power for the true complexity. We hereby formalize this as a testable core conjecture:
Conjecture 6.1 (Near-Tightness of RCC). For a given hardware capability $\Gamma_{R}$, there exist constants $a(\Gamma_{R})$ and $b(\Gamma_{R})$ that depend only on $\Gamma_{R}$, such that for a broad class of physically "well-structured" families of states (e.g., stabilizer states, graph states, or families of states reachable by low-rank, high-occupation projections), their optimal number of atomic instruction steps $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ satisfies the following upper bound:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (6.1)
where the constants $a$ and $b$ are assumed to be independent of the system dimension $d_{R}$ and the target state $\rho$ [^3] A detailed discussion of this parameterization is provided in Appendix G.. Validating or falsifying this conjecture would greatly deepen our understanding of the optimality of quantum circuit synthesis and could provide new information-theoretic guiding principles for the design of efficient quantum algorithms. Although proving this conjecture in a fully universal sense is challenging, we have found strong constructive support for its validity. Appendix G systematically proves that for three broad classes of "well-structured" state families prevalent in quantum information and physics, the upper bound on their optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ does indeed follow the form predicted by the conjecture. These three classes of states specifically include: (i) classical diagonal states whose amplitudes can be efficiently prepared; (ii) low-rank, high-occupation states that can be deterministically constructed through hierarchical projections; and (iii) efficiently encodable logical pure states with sparse generators. For these physically crucial states, the circuit complexity upper bounds we construct exhibit a consistent scaling behavior with $C_{R}(\rho)$, thus providing solid, non-trivial corroborating evidence for the near-tightness of Conjecture 6.1.
(Note: To precisely define the scope of this conjecture, it must be pointed out that it does not apply to Haar-typical pure states. For such states, the known optimal circuit complexity $C_{\mathrm{opt}}(\rho)$ satisfies an exponential scaling $\Omega(d_{R})$, whereas their Reference-Contingent Complexity $C_{R}(\rho)$ is merely $(\log d_{R})/\log\Gamma_{R}$. This exponential gap between $C_{\mathrm{opt}}$ and $RCC$ fundamentally restricts the domain of applicability of this conjecture to the "well-structured" state families with special constructions, as referred to in this paper. In particular, this qualitative gap also illustrates that for families of states that are "near-random and lack compressible structure," the RCC, as a lower bound on structured manufacturing cost, is intentionally "conservative," meaning it will not mistakenly judge "randomness" as "expensive structure.")
6.3 Resolving the Paradox of Pure State Complexity through an Observational Window
The core definition of the RCC framework, $C_{R}(\rho)=(\log d_{R}-S(\rho))/\log\Gamma_{R}$, while ensuring structural fairness, also introduces a profound paradox: from this global perspective, all pure states ($S(\rho)=0$)—whether they are simple product states or graph states with highly non-local entanglement—possess the same maximal leading term of complexity, $\log(d_{R})/\log(\Gamma_{R})$. This conclusion of "pure state degeneracy" severely contradicts both our physical intuition and practical computational experience, as it seems to completely erase the vastly different internal correlation structures of different quantum states, which are precisely what determine the true computational cost of preparing them. The resolution to this paradox is rooted in a more profound physical principle: complexity is not an intrinsic, absolute property of a quantum state, but rather a relational property that emerges from the interaction between an observer and the system. The structural complexity of a state is "latent" within its internal global correlations [41];to quantify it, one must actively "reveal" it through a physical probe. This "physical probe" is precisely the "observational window" $E_{\Xi}$ introduced for thermodynamic analysis in Sec. 5.4. This mathematical tool formalizes a physically realistic observer with finite capabilities. Such an observer cannot perceive the entire global pure state of the system at once, but can only interact with the system through a specific perspective limited by their capabilities (e.g., being able to perform only local measurements). The core insight is that when such a constrained observer examines a pure state with a complex internal structure, the state will appear as a mixed state within their "window" due to the inability to resolve its delicate global correlations, thus giving rise to an "emergent" non-zero entropy $S(E_{\Xi}(\rho))>0$. Crucially, the magnitude of this emergent entropy is a direct "fingerprint" of how the internal correlation structure of the pure state responds to that specific mode of observation. A window that is highly "aligned" with the state’s structure will reveal its latent complexity to the greatest extent; whereas a simple product state will remain simple under any local window, and its emergent entropy will be minimal. Therefore, by shifting the theoretical framework from the global $C_{R}(\rho)$ to the observation-window-dependent $C_{R,\Xi}(\rho)$, we not only solve the technical problem of pure state degeneracy but, more importantly, we elevate RCC into a more complete, dynamic theory. This theory acknowledges the central role of observation in defining physical reality [42] and provides a conceptually profound and practically feasible pathway for quantifying the structural complexity of different pure states. To this end, we further rigorize the concept of the observational window $E_{\Xi}$ from Sec. 5.4, defining it as a conditional expectation acting on $H_{R}$: $E_{\Xi}$ is a unital, completely positive and trace-preserving (CPTP), and idempotent quantum channel, corresponding to some von Neumann subalgebra $A_{\Xi}\subseteq\mathfrak{B}(H_{R})$, and satisfying the condition of compatibility with the reference vacuum, $E_{\Xi}(\sigma_{R})=\sigma_{R}$. This compatibility is an extension of the principle of structural fairness, ensuring that the observation process does not distort the zero point of complexity. Using the monotonicity of relative entropy, $D(\rho||\sigma_{R})\geq D(E_{\Xi}(\rho)||E_{\Xi}(\sigma_{R}))$, and the fact that $E_{\Xi}(\sigma_{R})=\sigma_{R}$, one can obtain $S(E_{\Xi}(\rho))\geq S(\rho)$, and thus $C_{R,\Xi}(\rho)\leq C_{R}(\rho)$. More importantly, this algebraic structure endows complexity with a hierarchical monotonicity: if the observational capability is enhanced, i.e., $A_{\Xi_{1}}\subseteq A_{\Xi_{2}}$, then the idempotency of the conditional expectation ensures $E_{\Xi_{1}}\circ E_{\Xi_{2}}=E_{\Xi_{1}}$, which in turn implies $S(E_{\Xi_{1}}(\rho))=S(E_{\Xi_{1}}(E_{\Xi_{2}}(\rho)))\geq S(E_{\Xi_{2}}(\rho))$, thereby yielding a well-behaved complexity hierarchy $C_{R,\Xi_{1}}(\rho)\leq C_{R,\Xi_{2}}(\rho)\leq C_{R}(\rho)$. To connect this theoretical framework with physical practice, we provide an operational lower bound, which forms the basis for experimentally probing the structural complexity of pure states.
Proposition 6.2 (Windowed Pinching Lower Bound). Let $\{P_{i}\}\subset A_{\Xi}$ be a set of orthogonal projection operators forming a complete set on $H_{R}$ ($\sum_{i}P_{i}=I_{H_{R}}$), with ranks $r_{i}=\mathrm{Tr}(P_{i})$. Let $p_{\Xi,i}=\mathrm{Tr}[P_{i}E_{\Xi}(\rho)]$ be the probability of observing outcome $i$ under the window $\Xi$. Then the windowed RCC (this paper uniformly uses the natural logarithm, with entropy in units of nats, consistent with the convention in Sec. 5.4 and Appendix I) satisfies the following inequality:
$C_{R,\Xi}(\rho)\geq\frac{\ln d_{R}-H(\{p_{\Xi,i}\})-\sum_{i}p_{\Xi,i}\ln r_{i}}{\ln\Gamma_{R}}$ (6.2)
In particular, when $\{P_{i}\}$ is a set of rank-one projectors (i.e., a basis measurement, in which case $S(P_{i}E_{\Xi}(\rho)P_{i}/p_{\Xi,i})=0$), this lower bound simplifies to:
$C_{R,\Xi}(\rho)\geq\frac{\ln d_{R}-H(\{p_{\Xi,i}\})}{\ln\Gamma_{R}}$ (6.3)
Proof Outline. Let $\sigma=E_{\Xi}(\rho)$. Apply a pinching map $\Delta_{P}(\sigma)=\sum_{i}P_{i}\sigma P_{i}$ to $\sigma$. By the entropy-increasing property of pinching, $S(\sigma)\leq S(\Delta_{P}(\sigma))$. At the same time, the entropy of the pinched state satisfies the block-entropy decomposition $S(\Delta_{P}(\sigma))=H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}S(\sigma_{i})\leq H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}\ln r_{i}$, where $\sigma_{i}=P_{i}\sigma P_{i}/p_{\Xi,i}$. Substituting this chain of inequalities back into the definition of $C_{R,\Xi}(\rho)$, $(\ln d_{R}-S(\sigma))/\ln\Gamma_{R}$, yields Eq. (6.2). (For technical details on the entropy increase due to pinching and the block-entropy decomposition used in this proposition, see Ref. [22]; for the rank-one dephasing entropy bound and experimental protocol, see D; for the axiomatic background of windowed monotonicity and notational/unit conventions, see Appendix I.) The physical insight of Eq. (6.2) is particularly crucial: it provides an operational "lever" to break the complexity degeneracy of pure states under a finite observational capability. For any pure state, the upper bound of the leading term of its ideal global complexity $C_{R}(\rho)$ is $\ln(d_{R})/\ln(\Gamma_{R})$. However, through a fixed local observational window $E_{\Xi}$, the value of $S(E_{\Xi}(\rho))$ for different pure states will differ significantly due to their different entanglement structures, thus creating a gap in their $C_{R,\Xi}(\rho)$ values. The significance of Eq. (6.3) is that it transforms $C_{R,\Xi}(\rho)$ from a theoretical construct into a physical quantity with direct operational meaning. While the definition of $C_{R,\Xi}(\rho)$ itself already conceptually solves the pure state degeneracy problem through $S(E_{\Xi}(\rho))$, Eq. (6.3) provides, from a more practical direction, an operational lower bound determined entirely by the classical measurement outcomes $\{p_{\Xi,i}\}$. This forms a key bridge connecting abstract theory with laboratory reality and gives rise to a powerful scheme for actively probing the latent complexity within pure states, which we term the "Experimental Squeeze of the Complexity Spectrum." Its core logic is that the choice of the measurement basis $\{P_{i}\}$ reflects the experimenter’s "informed probing" of the system’s intrinsic structural complexity. For a highly structured pure state, such as a graph state defined by a stabilizer group $S$, if we "intelligently" choose a measurement basis that is "aligned" with its stabilizer generators, then according to stabilizer theory, the measurement outcomes will be highly deterministic ($\pm 1$). This will cause the probability distribution $\{p_{\Xi,i}\}$ to become extremely sparse, with its Shannon entropy $H$ approaching zero. According to Eq. (6.3), this will cause its measurable complexity lower bound $C_{R,\Xi}(\rho)$ to be "squeezed" upwards, approaching its theoretical upper limit $\ln(d_{R})/\ln(\Gamma_{R})$. Conversely, a simple product state, lacking intrinsic non-local correlations, will inevitably appear more random when probed with the same complex measurement basis, resulting in a larger $H$ and a lower bound. This sharp contrast reveals a profound principle: the structural complexity of a state can only be most fully revealed when it is probed by an observer who "understands" its structural complexity. The significance of this "experimental squeeze" scheme is twofold. First, it constructs a feasible verification path for testing the core open question of this paper, Conjecture 6.1 (the near-tightness of RCC). By finding the optimal measurement bases that maximize the complexity lower bound for different classes of pure states (such as graph states, tensor network states, etc.), we can systematically establish "certificates of tightness" for their complexity. Second, it elevates the RCC framework from a static spectral measure to a complete, observation-dependent dynamical theoretical framework, in which the operational process of "finding the optimal measurement basis" is itself equivalent to the core task of "decoding the structural complexity of the quantum state."
6.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit
The core contribution of this work is the establishment of a rigorous information-theoretic lower bound (Theorem 3.1) for any quantum process that prepares a given state, derived from the structured information of the final state. This bound is essentially a static constraint, uniquely determined by the information content of the final state $\rho$ relative to its structured vacuum $\sigma_{R}$, and is independent of the specific evolution path. A natural question follows: does a complementary constraint exist, one that originates from the process dynamics itself, to limit the efficiency of the evolution? Quantum Speed Limits (QSLs) provide the theoretical framework for this. Recent advancements, particularly the "exact-QSL," have established an upper bound on this speed, determined by the variance of the non-classical part of the generator relative to the instantaneous state—a bound that is tighter than the Mandelstam–Tamm bound [43] [44]. In this section, we juxtapose the RCC information-theoretic lower bound with the exact-QSL dynamical upper bound to construct a dual-constraint theoretical framework for quantum processes, revealing the complementary relationship between information cost and dynamical efficiency. We first consider a pure-state unitary evolution process driven by a Hamiltonian $H$, evolving from an initial state $|\psi_{0}\rangle$ to a final state $\rho_{T}:=|\psi_{T}\rangle\langle\psi_{T}|$. Following Hall’s exact uncertainty relation [45], the generator can be decomposed relative to the instantaneous state $\rho_{t}$ as $H=H_{\text{cl}}+H_{\text{nc}}$, where only the uncertainty of the non-classical part, $\Delta H_{\text{nc}}$, contributes to the instantaneous speed of the quantum state ray [44]. The energy uncertainty of an operator $X$ is defined as $\Delta X(t)=\sqrt{\langle X^{2}\rangle_{t}-\langle X\rangle_{t}^{2}}$, and the time average is defined as $\langle f\rangle_{T}:=T^{-1}\int_{0}^{T}[f(t)]\,dt$.
Definition 6.4.1 (Fubini-Study Angle). For an initial state $|\psi_{0}\rangle$ and a final state $|\psi_{T}\rangle$, the Fubini-Study angle $\theta$ between them is defined as[44]:
$\theta=\arccos(|\langle\psi_{0}|\psi_{T}\rangle|)\in[0,\pi/2]$ (6.4)
$\theta$ quantifies the shortest geometric distance between the two quantum state rays. The geometric distance that any quantum evolution process can "travel" in the state space is strictly limited by the cumulative effect of the Hamiltonian’s "non-classical driving force" throughout the process.
Theorem 6.4.1 (exact-QSL, adapted from [44]). For any pure-state unitary evolution, its Fubini-Study angle $\theta$ is constrained by the following dynamical upper bound:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H(t)]\,dt$ (6.5)
For a general dimension, an exact expression for the evolution time $T$ is:
$T=\hbar\frac{l(\phi_{t})|_{0}^{T}}{\langle\Delta H_{\text{nc}}\rangle_{T}}$ (6.6)
where $l(\phi_{t})$ is the Wootters/statistical distance [46], and $\theta\leq l(\phi_{t})|_{0}^{T}$. In cases such as an effective two-level system where the survival probability $p_{t}=|\langle\psi_{0}|\psi_{t}\rangle|^{2}$ is monotonic, or when the Hamiltonian $H^{2}=I$ (self-inverse), the equality $\theta=(1/\hbar)\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ holds. [^4] For a single-qubit system with a self-inverse Hamiltonian $H=\sigma_{x}$ (taking $\hbar=1$) and initial state $|\psi_{0}\rangle=|0\rangle$, the evolved state is $|\psi_{t}\rangle=\cos(t)|0\rangle-i\sin(t)|1\rangle$. Here, $\theta(t)=t$ and $\Delta H_{\text{nc}}(t)=1$, so the equality in (6.5) holds, providing a minimal, achievable example of this theory. We can now juxtapose this dynamical ceiling with the RCC information-theoretic floor from this work. The two bounds apply to different objects (circuit cost vs. state geometric distance) and do not form a sandwich bound on the same quantity.
Information-Theoretic Floor:
$C_{\text{opt}}^{(\epsilon)}(\rho_{T})\geq\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots$ (6.7)
where $\dots"$ includes the universal logarithmic correction and the spectral skewness term $\Delta_{\text{spec}}$ (see Theorem 3.1); the dimension is normalized by $\log(\Gamma_{R})$ (in units of $st_{R}=\log_{2}(\Gamma_{R})$), consistent with the logarithmic base of $D(\cdot\|\cdot)$.
Dynamical Ceiling:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ (6.8)
This framework reveals a physical picture: the total computational cost ($C_{\text{opt}}$) of any successful quantum preparation process must be high enough to surpass the "information-theoretic threshold" set by the final state’s information structure, while, simultaneously, its evolutionary efficiency at any moment is constrained by a "dynamical speed limit" determined by the non-classical part of the Hamiltonian. To explore the connection between these two on a heuristic level, we can introduce a complexity proxy based on geometric distance.
Definition 6.4.2 (State Geometric Complexity). For a given initial state $|\psi_{0}\rangle$, we define the state geometric complexity $C_{\text{FS}}$ of the final state $\rho_{T}$ as its Fubini-Study angle:
$C_{\text{FS}}(\rho_{T})=\theta$ (6.9)
A state with higher information-structural complexity (large RCC) typically implies that it contains more complex quantum correlations and a more refined distribution of probability amplitudes. Starting from a simple initial state (such as a vacuum or product state), constructing these complex structures intuitively requires the quantum state to traverse a longer, non-trivial geometric path in Hilbert space. This physical intuition inspires us that a positive correlation may exist between the state’s information-theoretic complexity (RCC) and its geometric complexity (quantified by the proxy $C_{\text{FS}}=\theta$). We therefore propose the following heuristic alignment:
$\text{RCC}_{R}(\rho_{T})\lesssim C_{\text{FS}}(\rho_{T})=\theta$ (6.10)
The $\lesssim$ symbol in Eq.(6.10) emphasizes that this is a heuristic alignment, not a universal proposition. This relation suggests that a state with higher information-structural complexity (large RCC) may also need to traverse a greater geometric distance to be reached from a simple initial state. It is crucial to contextualize this heuristic. In other frameworks, such as Nielsen’s geometric complexity which employs a Finsler metric with penalty factors, the relationship is often reversed, yielding a model-dependent lower bound of the form $C_{\text{Nielsen}}\geq\lambda_{\gamma}\theta$ [47]. Our heuristic alignment, therefore, specifically pertains to the comparison between the information content (RCC) and the minimal geometric path ($\theta$). With this understanding, we adopt a conservative formulation for experimental purposes, proposing the hypothesis $\theta\geq c_{h}\,RCC_{R}(\rho_{T})$, where $0<c_{h}\leq 1$ is a model-dependent heuristic constant. Combining this conservative hypothesis with the traditional QSL bound $T\geq\hbar\theta/\langle\Delta H\rangle_{T}$ leads to a meaningful physical conclusion.
Corollary 6.4.1 (Auditable Minimum Time Bound, Heuristic). If there exists a heuristic constant $c_{h}\in(0,1]$ such that $\theta\geq c_{h}\cdot\text{RCC}_{R}(\rho_{T})$, then the evolution time $T$ is constrained by the following lower bound, determined by the final state’s information structure:
$T\geq\frac{\hbar}{\langle\Delta H\rangle_{T}}\cdot\theta\geq\frac{\hbar c_{h}}{\langle\Delta H\rangle_{T}}\cdot\left[\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots\right]$ (6.11)
This result connects the final state’s RCC information threshold with the process’s time-averaged energy uncertainty $\langle\Delta H\rangle_{T}$, collectively providing an auditable lower bound on the minimum evolution time. The dual-constraint framework proposed in this section unifies the static constraints of information theory and the dynamic constraints of dynamics, offering a more comprehensive and profound perspective for evaluating the resource consumption and evolutionary efficiency of quantum processes. It suggests that a deeper intrinsic unity may exist between computation, information, and dynamics than previously thought. Future open questions include: on which families of physically realizable systems does a stable quantitative relationship exist between RCC and $\theta$, and how to consistently generalize this framework to open quantum systems [48, 49].
6.5 Outlook for Future Work
The establishment of the RCC theoretical framework opens up avenues for future exploration in several directions. Firstly, the generalization from state complexity to the complexity of unitary operators/channels is a natural direction. By applying the idea of reference sets to Liouville space, we might be able to define an RCC for a unitary operator or a quantum channel, to quantify the intrinsic complexity of realizing a specific quantum operation compared to "free" operations (such as operations that preserve a specific subspace). Secondly, systematically connecting RCC with open quantum systems and quantum thermodynamics is another promising direction. For instance, by modeling processes that preserve energy eigen-populations or pinching processes of specific configurations as R-preserving channels, one can use the monotonicity of RCC, Eq. (5.1), to provide a non-trivial upper bound on the rate of information dissipation. Comparing this bound quantitatively with the change in thermodynamic quantities such as entropy production rate or non-equilibrium free energy is expected to lead to a series of testable inequalities connecting information-theoretic costs with thermodynamic prices. Finally, deepening the application of RCC in holographic duality will be an exciting long-term goal. As described in Sec. 5.3, RCC can serve as a structured calibration benchmark to test the order-preserving capability and proportional stability of different geometric candidate quantities (such as the CV/CA conjectures) across multiple families of states. If inconsistencies are found, they may suggest the existence of structural biases in the duality proposal and provide clues for finding directions for correction. (We provide a rigorous framework in Appendix F.2 for calculating a controllable, conservative lower bound on complexity in the presence of energy or mode cutoffs.)
7 Conclusion
This paper has systematically proposed and constructed the theoretical framework of “Reference-Contingent Complexity” ($RCC$), aiming to provide a self-consistent and operational solution for addressing the long-standing “structural fairness problem” in quantum information science. Starting from first principles, we formalize the intrinsic structure of a system—whether it originates from symmetries, conservation laws, or deliberate encoding—as an a priori “reference set,” and, using this as a benchmark, we define complexity as the information gap of a quantum state relative to this structured background, a gap which can be measured by relative entropy. This construction fundamentally ensures that our measure only prices the non-trivial quantum information that must be generated at the cost of computational resources. The core theoretical contribution of this paper is the proof of a main theorem (Theorem 3.1), which establishes that $RCC$ is a rigorous, theorem-level lower bound on the optimal number of atomic instruction steps for any universal quantum circuit. This theorem is not only mathematically rigorous, but its physical meaning also reveals a profound connection between the information-theoretic properties of a state (the information gap $D(\rho||\sigma_{R})$ and the spectral skew $\Delta_{\mathrm{spec}}$) and the computational cost of the dynamical process. To ensure the physical relevance of this theory, we have further proposed a set of operational principles based on quantum hypothesis testing, projective witnesses, and dephasing entropy bounds (detailed in Chapter 4), demonstrating that the lower bound from $RCC$ can be estimated through experiments and numerical simulations, thereby connecting an abstract theoretical concept with concretely measurable physical observables. As a theoretical tool, the value of $RCC$ is manifested in its broad application interfaces. We have shown the well-behaved mathematical properties it possesses as a proper resource measure, such as monotonicity under free operations (Sec. 5.1). We have elucidated how it complements and deepens the theoretical framework for the performance frontier of quantum dynamics, elevating a process-dependent complexity measure to a state function determined solely by the final state, as shown in (Sec. 5.2). Furthermore, we propose that the structural fairness of $RCC$ makes it a promising candidate to play a key role as a calibration benchmark in the exploration of fundamental physical questions such as the “complexity-geometry” holographic duality (Sec. 5.3). In summary, Reference-Contingent Complexity provides a foundational and structurally fair theoretical tool for understanding and quantifying the physical costs of quantum computation. It offers a unified “meter stick” for us to compare and diagnose increasingly complex quantum algorithms and hardware platforms, and opens up a solid and operational new path for exploring the deep unification of information, dynamics, and spacetime geometry. We are well aware that the establishment of any theoretical framework is merely the beginning of an exploration, not its end. Many theoretical properties of $RCC$, such as the tightness of its lower bound (Conjecture 6.1), its generalization to systems with non-Abelian symmetries, and its quantitative connection to quantum chaos and thermodynamics, remain open questions full of opportunities. We hope that this work will inspire more theoretical and experimental explorations within the field, jointly advancing towards a deeper understanding of the fundamental concept of “complexity” in the quantum world.
Appendix A Detailed Mathematical Proof of the Main Theorem
This appendix aims to provide a complete and mathematically rigorous proof for Theorem 3.1 in the main text. Our argument is divided into two core parts. First, in Section A.1, we provide a formal axiomatic definition of the quantum computation model, which serves as the cornerstone of the entire theoretical framework, with the goal of objectively mapping the computational cost of a physical process to an encoding length in an information-theoretic sense. Subsequently, in Sections A.2 to A.5, using this as a logical starting point, we systematically connect profound results from computational complexity, algorithmic information theory, and one-shot quantum information theory to construct, step by step, the complete inequality chain from the optimal number of atomic instruction steps $C_{\mathrm{opt}}$ to the Reference-Contingent Complexity $RCC$, thereby completing the proof of the main theorem.
Notations and Conventions. Within this appendix, all logarithms $\log$ are taken to base $2$; smoothing is performed using the trace distance $(1/2)\cdot||.||_{1}$. It is assumed by default that the support of the quantum state $\rho$ is contained within the subspace $H_{R}=\Pi_{R}\cdot H$; if physical leakage exists, the physical state can first be projected onto and normalized within $H_{R}$ before applying the conclusions that follow. The reference state is fixed as $\sigma_{R}=\Pi_{R}/d_{R}$, and the information bandwidth parameter is fixed as $\Gamma_{R}=g\cdot|S_{R}|$.
A.1 Formalization of the Computational Model
Our theoretical framework is established upon a discrete quantum computation model based on a universal gate set. To objectively quantify computational resources, we must precisely calibrate the information content of the fundamental unit of computation—the “atomic instruction”.
Definition A.1 (Alphabet of Atomic Instructions). An atomic instruction is an ordered pair $(u,s)$, where:
  1. 1.
    $u$ is a unitary operator selected from a universal gate set $G\subseteq\mathrm{SU}(d)$ with cardinality $g$.
  2. 2.
    $s$ is a target of action selected from a set of “addressable operational units” $S_{R}$ of size $|S_{R}|$.
The specific form of $S_{R}$ is determined by the physical connectivity of the hardware or the definition at the logical layer. Accordingly, the minimum number of information bits required to describe any single atomic instruction is determined by the total size of the alphabet of atomic instructions, $\Gamma_{R}$:
$\Gamma_{R}=g\cdot|S_{R}|$ (A.1)
Explanation (On Parallel Circuits and Serial Cost). In this paper, the optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ is directly defined as the minimum number of serial atomic instruction steps required to prepare the quantum state $\rho$ within a precision $\varepsilon$. For a physically parallel circuit, its depth and degree of parallelism jointly determine its equivalent total number of serial instructions. By directly defining $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ as the total serial cost, our encoding argument can be unambiguously benchmarked against it without needing to explicitly handle the details of compiling from parallel to serial.
Note on the Role of Continuous-Parameter Gates. The composite cost of continuous-parameter gates is already reflected in the optimal number of atomic instruction steps, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$; at the level of description length, the additional information written down for the target precision requires only $O(\log(1/\varepsilon))$ bits, and thus the corresponding overhead is incorporated into the logarithmic correction term.
(Note: Formalization of the computational model. For the purpose of the rigorous mathematical proofs that follow, this appendix adopts the computational model already established in the main text. The physical meaning and conventions for the optimal circuit complexity $C_{opt}$ and the information bandwidth parameter $\Gamma_{R}$ (particularly concerning how to handle parallel circuits and continuous-parameter gates) have been fully and unambiguously specified in the notes of Section 2.1 of the main text.)
A.2 Encoding Upper Bound: From Optimal Instruction Steps to Kolmogorov Complexity
Based on the formalized computational model in Section A.1, we can now encode any physical process that prepares a target quantum state $\rho$ into a descriptive program within the framework of algorithmic information theory. The universal machine $U_{R}$ takes the description of the reference set $R$ as a conditional input, and the term $K_{\varepsilon}(\rho|R)$ below is to be understood with this conditionality.
Lemma A.2 (Encoding Upper Bound). Let $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ be the minimum number of serial atomic instruction steps required to prepare the quantum state $\rho$ within a precision $\varepsilon$. Then, there exist universal constants $c_{a}$ and $c_{b}$ of order $O(1)$ such that the smoothed relative Kolmogorov complexity of $\rho$, $K_{\varepsilon}(\rho|R)$, satisfies the following inequality:
$K_{\varepsilon}(\rho|R)\leq C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}+c_{a}\cdot\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+c_{b}\cdot\log(1/\varepsilon)+O(1)$ (A.2)
[…]
(Prefix Code Specification. A self-describing prefix code (such as Elias-delta) is used for $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ and $\varepsilon$. Therefore, there exist constants $c_{a},c_{b}=O(1)$ that depend only on the chosen universal machine and the prefix code, and are independent of $\rho,\Gamma_{R},d_{R}$, such that Eq. (A.2) holds.)
A.3 Information-Theoretic Lower Bound on Quantum Kolmogorov Complexity
Lemma A.3 (Information-Theoretic Lower Bound). For any $\varepsilon>0$, there exists a constant $c_{0}=O(1)$, which depends only on the relative universal machine $U_{R}$ and is independent of $d_{R}$, such that [24][23] [26]:
$K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (A.3)
Proof.
  1. 1.
[…]
  2. 2.
    $\sigma$-padding and comparison of constant factors. Take any $\alpha\in(0,1)$ and define a new “padded” semi-density operator $M_{\alpha}=(1-\alpha)\cdot M+\alpha\cdot\sigma_{R}$. By construction, we have the evident operator orderings $\sigma_{R}\leq\alpha^{-1}\cdot M_{\alpha}$ and $M\leq(1-\alpha)^{-1}\cdot M_{\alpha}$. Using the scaling comparison property of $D_{\max}$: if $X\leq kY$, then $D_{\max}(.||Y)\leq D_{\max}(.||X)+\log k$. These relationships lead to: $D_{\max}(\omega||\sigma_{R})\leq D_{\max}(\omega||M_{\alpha})+\log(1/\alpha)$ and $D_{\max}(\omega||M_{\alpha})\leq D_{\max}(\omega||M)+\log(1/(1-\alpha))$. The inequalities remain valid after taking the infimum for $\varepsilon$-smoothing. Combining them yields $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||M)+\log(1/(\alpha(1-\alpha)))$. Since $M_{\alpha}\geq\alpha\cdot\sigma_{R}$, the support $\mathrm{supp}(M_{\alpha})$ contains $\mathrm{supp}(\sigma_{R})=H_{R}$, and thus all $D_{\max}(.||M_{\alpha})$ are finite and well-defined (without needing to normalize $M_{\alpha}$).
  3. 3.
    Choosing the constant and concluding. Combining the two steps above, we obtain $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq K_{\varepsilon}(\rho|R)+\log(1/(\alpha(1-\alpha)))$. To minimize the constant term, we choose $\alpha=1/2$, for which $\log(1/(\alpha(1-\alpha)))=\log(4)=2$ (with logarithm base 2). Hence: $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-2$ By incorporating a fixed $O(1)$ constant (denoted $c_{U}$) arising from the choice of the universal machine $U_{R}$ itself into the right-hand side, we can set $c_{0}=c_{U}+2$. This constant is clearly independent of $d_{R}$. Q.E.D.
A.4 Entropy Identity
Proposition A.4 (Entropy Identity). Under the premise that $\mathrm{supp}(\rho)\subseteq H_{R}$ and $\sigma_{R}=\Pi_{R}/d_{R}$,
$D_{\max}(\rho||\sigma_{R})-D(\rho||\sigma_{R})=S(\rho)-H_{\mathrm{min}}(\rho)=\Delta_{\mathrm{spec}}(\rho)$ (A.4)
Proof. On the subspace $H_{R}$, the operator $\log(\sigma_{R})$ is equivalent to $\log(d_{R})\cdot\Pi_{R}$. Since $\mathrm{supp}(\rho)\subseteq H_{R}$, we have $\mathrm{Tr}(\Pi_{R}\cdot\rho)=1$, and thus $\log(\sigma_{R})$ acts well on $H_{R}$. Therefore, the standard relative entropy is $D(\rho||\sigma_{R})=\mathrm{Tr}[\rho(\log(\rho)-\log(\sigma_{R}))]=\log(d_{R})-S(\rho)$. Furthermore, from the definition of max-relative entropy, the inequality $\rho\leq 2^{\lambda}\cdot\sigma_{R}$ is equivalent to $\rho||_{\mathrm{\infty}}\leq 2^{\lambda}/d_{R}$, which implies $D_{\max}(\rho||\sigma_{R})=\log(d_{R})-H_{\mathrm{min}}(\rho)$. Subtracting the two expressions yields the conclusion. Q.E.D.
A.5 Combining and Rearranging the Inequality Chain
We now chain together all the preceding results to complete the proof of Theorem 3.1. The core of the argument is to combine the information-theoretic upper bound on circuit cost (Lemma A.2) with the lower bounds on information-theoretic measures (Lemma A.3 and subsequent relations). First, we introduce a standard smoothing continuity bound. According to one-shot information theory, there exists a function $\alpha(\varepsilon)=O(\log(1/\varepsilon))$ that depends only on $\varepsilon$, such that:
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\geq D_{\max}(\rho||\sigma_{R})-\alpha(\varepsilon)$ (A.5)
Bounds of this type can be derived from the standard relationships between smoothed measures and the purified distance [50]. Second, by combining Lemma A.2 (Eq. (A.2)) and Lemma A.3 (Eq. (A.3)), we obtain the core inequality from physical cost to a smoothed information quantity. To ensure mathematical rigor, we uniformly use the form
$\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}):C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}+c_{a}\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+c_{b}\log(1/\varepsilon)+O(1)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (A.6)
Next, applying the aforementioned smoothing continuity bound (in the lower-bound direction) to the right-hand side and substituting the entropy identity proven in Proposition A.4 (Eq. (A.4)), we get:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}\geq D(\rho||\sigma_{R})+\Delta_{\mathrm{spec}}(\rho)-\alpha(\varepsilon)-c_{0}-c_{a}\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})-c_{b}\log(1/\varepsilon)-O(1)$ (A.7)
Finally, we divide both sides by $\log\Gamma_{R}$ and rearrange terms to obtain the final form. To consolidate all correction terms, we bound the universal constants: for any $\varepsilon\in(0,1/2]$, there exists a universal constant $\kappa=O(1)$ such that $\alpha(\varepsilon)\leq\kappa\cdot\log(1/\varepsilon)$. Based on this, we can choose a sufficiently large $O(1)$ constant $c_{1}$ such that it satisfies $c_{1}\geq\max\{c_{a},c_{b}+\kappa\}$, and also absorb the remaining $O(1)$ terms (such as $c_{0}$) into it. This yields the final form of the main theorem:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{D(\rho||\sigma_{R})}{\log\Gamma_{R}}-\frac{c_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}$ (A.8)
This expression is identical in form to Theorem 3.1 in the main text, where the coefficient of $\Delta_{\mathrm{spec}}(\rho)$ is exactly $1$, and all constants are independent of $d_{R}$. When an explicit non-recursive lower bound is needed, the solution method derived in Appendix C.2 can be applied to this bootstrap inequality.
A.6 Proof of the Smoothed Theorem and Tools from One-Shot Information Theory
This section aims to provide a self-contained and rigorous proof for Theorem 3.2 presented in the main text. This theorem connects the lower bound on circuit complexity with the hypothesis testing relative entropy $D_{H}^{\eta}$, which has a direct operational meaning. Our proof will be rooted in the modern framework of one-shot quantum information theory. This framework is specifically designed to handle information tasks in single, non-asymptotic systems. Its core tools are “smoothed” entropy quantities, which make the theoretical conclusions robust against the small perturbations that are unavoidable in physical systems.
A.6.1 Key One-Shot Information-Theoretic Quantities
We first precisely define the core information-theoretic quantities on which the proof in this section relies. By convention: all logarithms $\log$ are base-2, errors are measured by the trace distance $(1/2)\cdot||.||_{1}$, and it is assumed by default that the support of all quantum states $\rho$ is contained within the subspace $H_{R}$ defined by the reference set RR[26].
Definition A.6.1 (One-Shot Entropy Quantities).
  1. 1.
    Hypothesis testing relative entropy:
    $D_{H}^{\eta}(\rho||\sigma)=-\log\min_{0\leq T\leq I,\mathrm{Tr}[T\sigma]\leq\eta}\mathrm{Tr}[(I-T)\rho]$ (A.9)
    This quantity physically quantifies the minimum exponential decay rate of the type-II error probability when distinguishing between $\rho$ and $\sigma$ with an optimal strategy, under the constraint that the type-I error probability does not exceed $\eta$.
  2. 2.
    Max-relative entropy and its smoothed version:
    $D_{\max}(\rho||\sigma)=\inf\{\lambda:\rho\leq 2^{\lambda}\cdot\sigma\}$ (A.10)
[…]
A.6.2 Core Lemmas and Relations
Our proof relies on several standard results from one-shot information theory, which form the bridges between different entropy quantities.
Lemma A.6.1 (Data Processing Inequality). For any CPTP map $E$, the following entropy quantities are monotonically non-increasing:
$D_{H}^{\eta}(E(\rho)||E(\sigma))\leq D_{H}^{\eta}(\rho||\sigma)$ (A.12)
$D_{\max}^{\varepsilon}(E(\rho)||E(\sigma))\leq D_{\max}^{\varepsilon}(\rho||\sigma)$ (A.13)
Lemma A.6.2 (One-Shot Entropy Bridging Inequality). There exist system-independent universal constants $\kappa_{1},\kappa_{2}=O(1)$ such that:
$D_{H}^{\eta}(\rho||\sigma)\leq D_{\max}^{\varepsilon}(\rho||\sigma)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ (A.14)
Physical Interpretation. This inequality is the core bridge connecting the operational quantity $D_{H}^{\eta}$ with the more analytically tractable $D_{\max}^{\varepsilon}$. Its proof is rooted in the quantum Neyman-Pearson lemma and its smoothing techniques.
A.6.3 Upper Bound on Smoothed Max-Relative Entropy from Circuit Complexity
The first step of the proof is to establish a connection between the optimal number of circuit steps $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ and $D_{\max}^{\varepsilon}(\rho||\sigma_{R})$, which is in the same vein as the logic in Section A.2 but employs a more self-contained proof technique.
Lemma A.6.3 (Description Length Upper Bound). Let $L=C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$. There exist universal constants $a,b=O(1)$ that are independent of $d_{R}$, such that:
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)$ (A.15)
Proof. Following the prefix-free encoding idea from Section A.2, we construct the universal semi-density operator $M=\sum_{p}2^{-|p|}\rho_{p}$. If there exists a circuit described by a program $p^{*}$ that can $\varepsilon$-approximately prepare $\rho$ (with output $\rho_{p^{*}}$), then by the definition of $M$, we have the operator inequality $M\geq 2^{-|p^{*}|}\cdot\rho_{p^{*}}$, which directly leads to $D_{\max}(\rho_{p^{*}}||M)\leq|p^{*}|$. To change the comparison object from $M$ to $\sigma_{R}$, we introduce a mixed semi-density operator $M_{\alpha}=(1-\alpha)M+\alpha\sigma_{R}$, where $\alpha\in(0,1)$. Using the scaling comparison property of max-relative entropy, we get $D_{\max}(\rho_{p^{*}}||\sigma_{R})\leq D_{\max}(\rho_{p^{*}}||M)+\log(1/(\alpha(1-\alpha)))$. Choosing $\alpha=1/2$ makes the constant term $\log(1/(\alpha(1-\alpha)))=\log(4)=2$. Combining this with the information-theoretic upper bound on $|p^{*}|$, which is $L\cdot\log\Gamma_{R}+O(\log L)$, and the continuity of smoothed entropies, we obtain Eq. (A.15). This proof path, via the $\sigma$-padding technique, avoids any spectral constants with explicit dependence on $d_{R}$.
A.6.4 Proof of Theorem 3.2
We now have all the elements necessary to prove the smoothed theorem. Theorem 3.2 (Restated). Under the same conditions as in Theorem 3.1, there exist system-independent universal constants $c^{\prime}_{1},c^{\prime}_{2}=O(1)$ such that for any $\eta,\varepsilon\in(0,1)$:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D_{H}^{\eta}(\rho||\sigma_{R})-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\log(1/\eta)$ (A.16)
Proof. Let $L=C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$. We chain together Lemma A.6.2 and Lemma A.6.3. From Eq. (A.14) and Eq. (A.15), we have: $D_{H}^{\eta}(\rho||\sigma_{R})\leq\left(L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)\right)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ By algebraically rearranging the above expression to solve for a lower bound on $L$, and defining the universal constants $c^{\prime}_{1}=\max\{a,b+\kappa_{2}\}$ and $c^{\prime}_{2}=\kappa_{1}$, we obtain Eq. (A.16).
A.6.5 Non-Bootstrap Closed-Form Lower Bound
Eq. (A.16) is a bootstrap-type inequality. For practical applications, we can provide a non-recursive, closed-form lower bound for it using the method derived in Appendix C.2. Let:
$A$ $=\frac{D_{H}^{\eta}\!\bigl(\rho\,\|\,\sigma_{R}\bigr)}{\log\Gamma_{R}},$ (A.17)
$B$ $=\frac{1}{\log\Gamma_{R}}\!\left(c^{\prime}_{1}\log\!\frac{1}{\varepsilon}\;+\;c^{\prime}_{2}\log\!\frac{1}{\eta}\right),$ (A.18)
$c$ $=\frac{c^{\prime}_{1}}{\log\Gamma_{R}}$ (A.19)
Then Eq. (A.16) can be written as $L\geq A-B-c\cdot\log(\max\{1,L\})$. Its smallest solution $L^{*}$ is given by the Lambert-W function $W_{0}$:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\max\left\{1,\left(\frac{c}{\ln 2}\right)\cdot W_{0}\left(\left(\frac{\ln 2}{c}\right)\cdot\exp\left(\left(\frac{\ln 2}{c}\right)\cdot(A-B)\right)\right)\right\}$ (A.20)
In scenarios where an implementation without special function libraries is required, the conservative approximate bound given in Appendix C.2 can be used.
A.6.6 Framework Consistency and Independence Check
Independence of Constants from $d_{R}$: All universal constants in the proof of this section originate from the self-description overhead of prefix codes and the bridging inequalities for one-shot entropies. By employing the proof technique involving the mixture $M_{\alpha}$, we avoid any spectral constants with explicit dependence on $d_{R}$, thereby ensuring the universality of the final constants $c^{\prime}_{1},c^{\prime}_{2}$. Alignment with Theorem 3.1: The present framework is fully self-consistent with the proof framework of Sections A.1-A.4. If this result is combined with the proven entropy identity (Proposition A.4) and standard entropy relations, one can recover the "three-part structure" of Theorem 3.1: $\text{leading term }D/\log\Gamma_{R}+\text{spectral correction }\Delta_{\mathrm{spec}}/\log\Gamma_{R}+\text{logarithmic correction}$, with the coefficients of the leading term and the spectral correction term both being exactly $1$.
Accounting for Errors and Compilation: The $O(\log(1/\varepsilon))$ corrections arising from smoothing and gate synthesis have been uniformly incorporated into the logarithmic correction term of Eq. (A.16); the constant factor from compiling parallel to serial circuits does not affect the scaling behavior and only influences the specific numerical values of the $O(1)$ coefficients $a$ and $b$.
Appendix B Technical Foundations and Application Context
This appendix aims to provide a series of key technical clarifications and supplementary materials for both theoretical and experimental researchers. Section B.1 elaborates on how this theoretical framework robustly handles the unavoidable problem of state “leakage” in physical systems. Section B.2 provides a key technical proof, ensuring the universality of the correction terms in the main theorem. Finally, Section B.3 places Reference-Contingent Complexity (RCC) within a broader theoretical landscape, systematically analyzing its relationship with other mainstream complexity paradigms, and proposes an operational protocol for lower-bound estimation.
B.1 Robustness of the Theory to Support Set Leakage
The theoretical construction in the main text, particularly the main theorem, Theorem 3.1, under ideal conditions assumes that the support of the target state $\rho$ lies strictly within the subspace $H_{R}$ defined by the reference set $R$, i.e., $\mathrm{supp}(\rho)\subseteq H_{R}$. However, in any physical experiment or numerical simulation, due to finite control precision or decoherence effects, the actually prepared state $\rho_{\mathrm{phys}}$ will always unavoidably have a small component that leaks into the orthogonal complement space $H_{R}^{\perp}$. This leakage results in $\mathrm{Tr}(\Pi_{R}\cdot\rho_{\mathrm{phys}})=1-p_{\mathrm{leak}}<1$. If such a $\rho_{\mathrm{phys}}$ were to be directly substituted into the definition of relative entropy $D(\rho_{\mathrm{phys}}||\sigma_{R})$, because the support of $\sigma_{R}$ on $H_{R}^{\perp}$ is null (its operator is the zero operator on this complement space), the relative entropy would strictly diverge. To ensure the applicability of this theoretical framework in realistic scenarios, we provide the following two standard schemes for handling this issue.
Scheme One: State Projection and Renormalization
The most direct way to handle this is to first project the physical state back into the target subspace and then proceed with the analysis.
Proposition B.1 (Construction of the Effective State): Given a physical state $\rho_{\mathrm{phys}}$, we can define its effective state $\tilde{\rho}$ within the structured subspace $H_{R}$ as:
$\tilde{\rho}=\frac{\Pi_{R}\cdot\rho_{\mathrm{phys}}\cdot\Pi_{R}}{\mathrm{Tr}(\Pi_{R}\cdot\rho_{\mathrm{phys}})}$ (B.1)
All conclusions in the main theorem, Theorem 3.1, can be directly applied to this effective state $\tilde{\rho}$. The physical meaning of this scheme is that we focus our analysis on the part of the quantum state that was successfully prepared within the target structure, treating it as the true computational goal. (Note: The above construction requires $q=\mathrm{Tr}(\Pi_{R}\rho_{\mathrm{phys}})>0$. When $q=0$ (complete leakage), the smoothed reference state scheme (B.2) should be used instead.)
Scheme Two: Smoothing of the Reference State
Another, more general method in information theory is to construct a slightly perturbed “smoothed” reference state that has support over the entire space.
Proposition B.2 (Construction of the Smoothed Reference State): For any small $\delta>0$, we can define a $\delta$-smoothed reference state $\tilde{\sigma}_{R}$ as:
$\tilde{\sigma}_{R}=(1-\delta)\cdot\sigma_{R}+\delta\cdot\frac{(I-\Pi_{R})}{(d-d_{R})}$ (B.2)
where $I$ is the identity operator on the total Hilbert space $H$. The relative entropy $D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})$ is now finite, and it can be shown that it provides a controllable approximation for $D(\tilde{\rho}||\sigma_{R})$. Their difference can be bounded by logarithmic terms related to the leakage probability $p_{\mathrm{leak}}$ and the smoothing parameter $\delta$. In practical applications, as long as the leakage probability $p_{\mathrm{leak}}$ is much smaller than the target precision $\varepsilon$, the corrections to the leading term of the complexity lower bound resulting from either scheme are secondary. Therefore, this theoretical framework is robust against the small state leakage that is physically unavoidable. (Note: Here $I$ is the identity operator on $H$, and $d=\dim H$. If $d_{R}=d$, then $\tilde{\sigma}_{R}=\sigma_{R}$, and no smoothing is necessary in this case.)
B.2 Proof of the $d_{R}$-Independence of the Kolmogorov Lower Bound Constant $c_{0}$
The universality of the main theorem, Theorem 3.1, relies on the constant $c_{0}=O(1)$ in the inequality $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (which was used as a lemma in Appendix A.3) not varying with the subspace dimension $d_{R}$. This section provides an independent, constructive proof for this key property, explicitly showing that $c_{0}$ can be taken as a constant independent of $d_{R}$.
Proposition B.3 ($d_{R}$-Independence of $c_{0}$). There exists a constant $c_{0}=O(1)$, whose magnitude depends only on the choice of the relative universal machine $U_{R}$, such that for any $\varepsilon>0$ and any $d_{R}$, it holds that $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$[51].
Proof. Let the universal semi-density operator for the relative universal machine $U_{R}$ be $M=\sum_{p}2^{-l(p)}\cdot\rho_{p}$, where $l(p)$ is the length of the shortest prefix code for program $p$, and $\rho_{p}$ is its output state, which by construction is guaranteed to have $\mathrm{supp}(\rho_{p})\subseteq H_{R}$. Therefore, $\mathrm{Tr}(M)\leq 1$ and $\mathrm{supp}(M)\subseteq H_{R}$.
  1. 1.
[…]
    $D_{\max}^{\varepsilon}(\rho||M)\leq L=K_{\varepsilon}(\rho|R)$ (B.3)
    (Note: This paper adopts the common extended definition from the one-shot literature, which allows $Y$ to be positive semi-definite with $\mathrm{Tr}Y\leq 1$; this differs from the normalized version by only an $O(1)$ constant, which has been absorbed into $c_{0}$.)
  2. 2.
    $\sigma$-padding and Scale Comparison: For any fixed $\alpha\in(0,1)$, define $M_{\alpha}=(1-\alpha)\cdot M+\alpha\cdot\sigma_{R}$. It is clear that $M_{\alpha}\geq\alpha\cdot\sigma_{R}$ and $M_{\alpha}\geq(1-\alpha)\cdot M$. Using the scaling comparison property of $D_{\max}$ (if $X\leq kY$, then $D_{\max}(.||Y)\leq D_{\max}(.||X)+\log k$), for any state $\omega$, we have:
    $D_{\max}(\omega||\sigma_{R})\leq D_{\max}(\omega||M_{\alpha})+\log(1/\alpha)$
    $D_{\max}(\omega||M_{\alpha})\leq D_{\max}(\omega||M)+\log(1/(1-\alpha))$
    The inequalities remain valid after taking the infimum for $\varepsilon$-smoothing, thus:
    $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||M)+\log(1/(\alpha(1-\alpha)))$ (B.4)
  3. 3.
    Combining and Parameter Selection: From Eq. (B.3) and Eq. (B.4), we have:
    $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq K_{\varepsilon}(\rho|R)+\log(1/(\alpha(1-\alpha)))$
    By choosing $\alpha=1/2$, we get $\log(1/(\alpha(1-\alpha)))=\log(4)=2$ (with logarithm base 2). Hence:
    $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-2$ (B.5)
    By further incorporating a fixed $O(1)$ constant (denoted as $c_{U}$) arising from the choice of the universal machine $U_{R}$ itself into the right-hand side, we can set $c_{0}=c_{U}+2$. This constant is clearly independent of $d_{R}$. Q.E.D.
(Note: We do not rely on any spectral comparison constants between $M$ and $\sigma_{R}$; using $\sigma$-padding (with $\alpha=1/2$) is sufficient to obtain $K_{\varepsilon}(\rho\!\mid\!R)\geq D_{\max}^{\varepsilon}(\rho\|\sigma_{R})-c_{0}$, where $c_{0}=O(1)$ and is independent of $d_{R}$, see Appendix A.3 and Section B.2.) [51]
B.3 Comparison with and Positioning Relative to Other Complexity Measures
This section aims to place Reference-Contingent Complexity (RCC) within the broader landscape of existing theories. Through a multi-dimensional comparison, it clarifies its unique theoretical positioning, application scenarios, and complementary relationship with other mainstream paradigms. The core contribution of RCC is that it provides a structurally fair, theorem-level lower bound for the complexity of any quantum circuit based on a universal gate set, and connects this theoretical concept with a set of operational measurement protocols.
Table 1: Table B.1: A Multi-dimensional Comparison of RCC with Other Complexity Measures
Dimension Reference-Contingent Complexity (RCC) Nielsen’s Geometric Complexity Krylov Complexity Resource-Theoretic Relative Entropy Mana/T-count
Object of Measure State function (final state) Path function (dynamical process) Operator growth (dynamics) Resource content of a state Resource count for a specific gate set
Structural Fairness Explicitly modeled via reference set Depends on the choice of penalty factors Partial (depends on basis choice) Yes (depends on the definition of the free set) No (model is fixed)
Relation to Gate Set Explicitly dependent via $\Gamma_{R}$ Implicitly dependent via metric Typically independent of gate set Related to the set of free operations Strongly dependent on a specific gate set
Bound on Circuit Complexity Theorem-level lower bound Typically an upper bound or an approximation Indirectly related Not a direct bound Strict lower bound for a specific model
Operationality Yes (protocol in Sec. B.3.2) Indirect (requires path optimization) Yes (measurable via response functions) Partial (depends on free operations) Yes (can be counted directly by compiling)
Path Dependence No Yes Yes No No
B.3.1 Analysis of Applicability in Specific Physical Scenarios
Different complexity paradigms have their respective focuses in different physical problems. The unique advantages of RCC lie in its structural fairness and its property as a universal lower bound[40, 38, 39].
  1. 1.
    Quantum Error Correction and Logical Computation: When assessing the preparation cost of logical states for different quantum error-correcting codes, a core challenge is to fairly compare different physical encoding schemes [14, 15]. By defining the code space as the reference set, RCC’s calculation result depends only on the complexity of the logical information and is independent of the number of physical qubits [52]. This makes it an ideal, encoding-independent performance benchmark [16]. It can be combined with model-specific resource counts (such as T-count for the Clifford+T model) to provide complementary perspectives [53, 54]. In addition, practical architectures illustrate how logical operations exploit code structure without inflating physical-qubit–level costs [17].
  2. 2.
    Hamiltonian Dynamics and Quantum Chaos: When studying the process of state complexification under Hamiltonian evolution, Krylov complexity—through the spreading of an operator in the Krylov basis—provides deep insights into the chaotic nature of the dynamics [9]. RCC can serve as a complementary tool: by selecting a sector of the system’s energy eigenstates as the reference set, one can study the minimum information-theoretic complexity that the final state of the system can achieve under energy constraints, thereby providing a lower bound on the range of states accessible to the dynamical process.
  3. 3.
    Quantum Resource Theories: Resource theories, such as those for “mana” or entanglement, often quantify the resource content of a state via relative entropy with respect to a set of free states, $D(\rho\|F)$ [55, 56]. RCC can be viewed as both a special case and a generalization: it concretizes “free” into a reference state $\sigma_{R}$ defined by an arbitrary physical structure and links it directly to computational cost [13]. For stabilizer-based nonclassicality and related resource pricing, representative instances include magic-state measures and operational characterizations [10, 11].
B.3.2 An Operational Protocol for Estimating the Complexity Lower Bound
To apply RCC in experiments or numerical simulations, we propose a lower-bound estimation protocol consisting of three independent paths, complete with statistical confidence.
  1. 1.
    Determine the Reference Set: First, determine the reference set $R$ based on the physical constraints of the system (e.g., stabilizer group, conserved quantity sector, or logical code space), and from this, calculate the subspace dimension $d_{R}$ and the reference state $\sigma_{R}$. Concurrently, calibrate the information bandwidth parameter $\Gamma_{R}$ based on the gate set and hardware architecture used.
  2. 2.
    Execute a Measurement Path: Choose at least one of the following paths for measurement to obtain a lower bound on the leading term of RCC, $D(\rho||\sigma_{R})$:
    1. (a)
      Path One (Quantum Hypothesis Testing): Frame the task of distinguishing $\rho$ and $\sigma_{R}$ as a binary hypothesis testing problem. Through repeated experiments, estimate the upper confidence endpoint $\beta_{U}$ of the type-II error probability (mistaking $\rho$ for $\sigma_{R}$). This yields a lower bound on $D_{H}^{\eta}(\rho||\sigma_{R})$, namely $\log(\beta_{U})$.
    2. (b)
      Path Two (Projective Witnesses): Select one or more pre-calibrated low-rank (rank $r$) projection operators $\Pi$ whose support is within $H_{R}$. Through measurement experiments, estimate the lower confidence endpoint $p_{\mathrm{lower}}$ of its occupation probability on the state $\rho$. This yields a lower bound on $D_{\max}(\rho||\sigma_{R})$, namely $\max\{0,\log(p_{\mathrm{lower}}\cdot d_{R}/r)\}$.
    3. (c)
      Path Three (Dephasing Information Bound): Perform measurements in a classical basis compatible with the reference set $R$ to reconstruct the probability distribution of the measurement outcomes. Estimate the upper confidence interval for the Shannon entropy $H$ of this classical distribution using statistical methods. This yields a lower bound on $D(\rho||\sigma_{R})$, namely $\log(d_{R})-H$.
  3. 3.
    Synthesize and Report: Take the maximum of the lower bounds obtained from the multiple paths above. Combine this with the spectral skew term and the logarithmic correction terms from the main theorem, Theorem 3.1, to finally report a lower bound on the optimal circuit complexity $C_{\mathrm{opt}}$, complete with explicit confidence levels and hardware assumptions. A comprehensive analysis should also compare this lower bound with upper bounds derived from Nielsen geometry or specific algorithm compilation, thereby forming a "sandwich" estimate that brackets the true complexity from above and below.
Appendix C Formal Details and Implementation Protocols
This appendix aims to consolidate and formally articulate several key technical details and implementation protocols mentioned in the main text. Its purpose is to provide a deeper level of mathematical support for the rigor of the theory and to offer a precise, operational guide for researchers who wish to apply this framework in experiments or numerical simulations[57, 58].
Base and Shorthand. In this appendix, all logarithms $\log$ are by default base $2$, and $\ln$ denotes the natural logarithm. For the sake of conciseness, let $c=c_{1}/\log\Gamma_{R}>0$, and $D=(D(\rho||\sigma_{R})+\Delta_{\mathrm{spec}}(\rho))/\log\Gamma_{R}-c\cdot\log(1/\varepsilon)$. The main inequality, Eq. (3.1), can be written compactly as $x\geq D-c\log(x)$, and its corresponding equation is $x+c\log(x)=D$.
C.1 A Physics Commentary on the Definition of $C_{\mathrm{opt}}$
To ensure the rigor of our theoretical derivations, in Appendix A, we directly define the optimal circuit complexity $C_{\mathrm{opt}}$ as the minimum required number of serial atomic instruction steps. This definition is clear and unambiguous within the framework of algorithmic information theory. Here, we aim to provide a supplementary explanation from a physics perspective for this formal definition, to clarify its connection with the parallel quantum circuit models widely used in physical reality. In concrete hardware implementations, quantum circuits are typically executed in a parallel fashion. The computational resources of a parallel circuit are usually characterized by two dimensions: the depth of the circuit, $D$ (i.e., the number of parallel time layers), and the degree of parallelism of each layer (i.e., the number of non-conflicting operations that can be performed simultaneously in that layer). The total computational workload, or total cost, of a circuit is physically proportional to the sum of all atomic operations performed across all parallel layers. Our definition of $C_{\mathrm{opt}}$ is, conceptually, a direct quantification of this total physical cost. It is equivalent to the total length of an optimal parallel circuit after it has been "flattened" into an equivalent serial instruction sequence. This method of definition not only maintains a close connection to physical reality but also maximizes the simplicity and universality of the theoretical derivations.
C.2 Closed-Form Solution and Usable Approximations for the Bootstrap Inequality
The main theorem, Eq. (3.1), leads to a bootstrap-type lower bound $x\geq D-c\log(x)$ (where $x>0$). This section provides its rigorous closed-form minimum solution, its monotonicity and uniqueness, and offers two types of implementable approximations for direct application in numerical or experimental reporting.
Lemma C.1 (Monotonicity and Unique Root). Define the function $F(x)=x+c\log(x)-D$. For all $x>0$, its derivative is $F^{\prime}(x)=1+c/(x\cdot\ln 2)>0$. Therefore, $F(x)$ is strictly monotonically increasing on $(0,\infty)$, and the equation $F(x)=0$ has a unique solution on $(0,\infty)$, denoted as $x_{*}$. This unique solution $x_{*}$ is the minimum possible value that satisfies the original inequality.
Theorem C.2 (Tight Closed-Form Solution via Lambert-W). Let $\alpha=c/\ln 2>0$. Then the minimum solution $x_{*}$ of the bootstrap inequality is given exactly by the principal branch $W_{0}$ of the Lambert-W function:
$x_{*}=\alpha\cdot W_{0}\left(\frac{1}{\alpha}\cdot\exp\left(\frac{D}{\alpha}\right)\right)$ (C.1)
Proof. The equation $x+c\log(x)=D$ is equivalent to $x+\alpha\ln(x)=D$. Rearranging gives $(x/\alpha)\cdot\exp(x/\alpha)=(1/\alpha)\cdot\exp(D/\alpha)$. According to the definition of the Lambert-W function, $u\exp(u)=z\iff u=W(z)$, by setting $u=x/\alpha$ and $z=(1/\alpha)\cdot\exp(D/\alpha)$, we can solve for $x/\alpha=W_{0}(z)$, where the principal branch $W_{0}$ ensures the unique positive real solution. Q.E.D.
Proposition C.3 (Practical Approximation I: Leading Term for Large $D$). When $D/\alpha$ is sufficiently large, using the asymptotic expansion of $W_{0}(z)$, which is $\ln(z)-\ln(\ln(z))+o(1)$, one can obtain the leading-term approximation of the minimum solution:
$x_{*}=D-c\log(D)+o(1)\quad\text{as }D\to\infty$ (C.2)
Proposition C.4 (Practical Approximation II: Piecewise Conservative Lower Bound). For any $D\in\mathbb{R}$ and $\alpha>0$, the following piecewise function provides a strict, conservative lower bound for $x_{*}$ that can be computed without requiring a special function library:
$x_{*}\geq\begin{cases}\max\{0,D/(1+\alpha)\}&\text{for }0<D\leq 1+\alpha\\ D-\alpha\ln(1+D)&\text{for }D>1+\alpha\end{cases}$ (C.3)
Implementation Recommendation: When performing numerical evaluations, the Lambert-W closed-form solution from Eq. (C.1) should be used preferentially. If a Lambert-W function library is not available, Eq. (C.2) can be used as an initial value for Newton’s method, or the conservative piecewise lower bound given by Eq. (C.3) can be reported directly.
C.3 Robustness to Support Set Leakage: Error Propagation
Lemma C.5 (Leakage Error Bound). There exists a universal constant $k=O(1)$ such that:
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D(\tilde{\rho}||\sigma_{R})-k\cdot\left(\log\left(\frac{1}{1-p_{\mathrm{leak}}}\right)+\log\left(\frac{1}{\delta}\right)\right)$ (C.4)
Proof Sketch. Let $\rho_{\mathrm{phys}}=q\tilde{\rho}\oplus(1-q)\rho_{\perp}$, where $q=1-p_{\mathrm{leak}}$. Take the smoothed reference state $\tilde{\sigma}_{R}=(1-\delta)\sigma_{R}\oplus\delta\tau_{\perp}$. Consider a binary measure-and-prepare channel $M$ for "whether the state falls within $H_{R}$". According to the data processing inequality and the chain rule for relative entropy, we have $D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq D_{\mathrm{Bern}}(q||1-\delta)+qD(\tilde{\rho}||\sigma_{R})$, where $D_{\mathrm{Bern}}$ is the binary classical KL divergence. Using the continuity bound for $D_{\mathrm{Bern}}$, the first term can be lower-bounded by a term of the form $k(\log(1/q)+\log(1/\delta))$, where $k=O(1)$ is a universal constant. From this, Eq. (C.4) follows.
(Note and Implementation: Choosing the smoothing parameter $\delta=p_{\mathrm{leak}}$ can minimize the $D_{\mathrm{Bern}}$ term, in which case the error bound can be simplified to a multiplicative reduction form $D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R})\geq(1-p_{\mathrm{leak}})\cdot D(\tilde{\rho}||\sigma_{R})$.)
C.4 Standard Relations for One-Shot Entropies
Proposition C.6 (Entropy Relations). There exist model-independent universal constants $k_{1},k_{2}=O(1)$ such that:
$D_{H}^{\eta}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||\sigma_{R})+k_{1}\cdot(\log(1/\eta)+\log(1/\varepsilon))$ (C.5)
$D(\rho||\sigma_{R})\geq D_{H}^{\eta}(\rho||\sigma_{R})-k_{2}\cdot\log(1/\eta)$ (C.6)
where all $\log$ are base $2$. These inequalities serve as the bridge connecting Theorem 3.1 and Theorem 3.2, and their proofs rely on standard techniques from one-shot quantum information theory [59].
C.5 Example Calculations and Implementation Guide
We first demonstrate the calculation method of RCC and its embodiment of "structural fairness" through two idealized physical systems.
  1. 1.
    Example 1 (Logical Pure State of a Stabilizer Code). For $k$ logical qubits encoded into $N$ physical qubits, the dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For any logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assume that logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and a single operation can address all $k$ logical qubits. Its RCC is:
    $C_{R}(\rho_{L})=\frac{k}{\log(g_{L}\cdot k)}$ (C.7)
    Implementation Recommendation: When reporting $C_{R}$, the information bandwidth $\Gamma_{R}=g_{L}\cdot k$ should be fixed. One can show a plot of $C_{R}$ as a function of increasing $k$, accompanied by comparative data showing that $C_{R}$ remains constant for different numbers of physical qubits $N$, to visually demonstrate "structural fairness".
  2. 2.
    Example 2 (Fixed Particle Number Sector). In an $n$-qubit Hilbert space, consider the subspace defined by the conservation of total particle number (Hamming weight $w$). This reference set $R$ is defined by a single projector $\Pi_{R}$, and its dimension is $d_{R}=\binom{n}{w}$, where $\binom{n}{w}$ is the binomial coefficient. Measure-and-Estimate Path: Perform measurements in a computational basis compatible with the $U(1)$ symmetry (such as the particle number basis) to obtain an empirical probability distribution $P_{\mathrm{hat}}$. Use the methods based on the DKW inequality and the Fannes-Audenaert continuity bound from Appendix D.3 to construct an upper confidence endpoint $H_{U}(\delta)$ for the Shannon entropy $H$. Finally, report a complexity lower bound with statistical confidence:
    $C_{R}\geq\frac{(\log d_{R}-H_{U}(\delta))}{\log\Gamma_{R}}$
    Beyond the idealized examples above, we provide a more universally applicable implementation guide. In an experiment, even if it is not possible to obtain full information about the state $\rho$, as long as there is any executable measurement (described by some POVM), we can extract a meaningful complexity lower bound. Based on a conservative lower bound from an arbitrary measurement, our recommended implementation guide is as follows: if only measurement sample data from a specific POVM is available, a conservative but effective method is to first calculate the classical KL divergence $D_{\mathrm{meas}}$ between the classical probability distributions induced by this POVM. This quantity is a rigorous lower bound for $D(\rho||\sigma_{R})$. Consequently, one can report a conservative lower bound on complexity $C_{R}(\rho)\geq\frac{(\log d_{R}-H_{\mathrm{meas}})}{\log\Gamma_{R}}$, where $H_{\mathrm{meas}}$ is the measurement entropy of $\rho$ under this POVM.
Appendix D Technical Details and Statistical Protocols for the Operational Principles
This appendix provides the technical details and statistical protocols for the three information extraction principles proposed in Chapter 4 of the main text: quantum hypothesis testing, projective witnesses, and the dephasing entropy bound. Its core objective is to provide testable lower bounds for $D_{H}^{\eta}(\rho||\sigma_{R})$, $D_{\max}(\rho||\sigma_{R})$, or $D(\rho||\sigma_{R})$ with rigorous statistical control at a confidence level of $1-\delta$, so as to facilitate the application of the main theorem to real data. We first unify the notation and prerequisites[57]:
  1. 1.
    In this appendix, $\log$ denotes the logarithm to base $2$; $\ln$ denotes the natural logarithm.
  2. 2.
    The reference state is $\sigma_{R}=\Pi_{R}/d_{R}$, where $d_{R}=\mathrm{Tr}(\Pi_{R})$.
  3. 3.
    Unless otherwise specified, it is assumed by default that $\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\Pi_{R})$. If leakage exists, one may perform projection and renormalization according to Appendix B.1 or use a smoothed reference state.
  4. 4.
    The gate bandwidth parameter is $\Gamma_{R}=g\cdot|S_{R}|\geq 2$, and therefore $\log\Gamma_{R}>0$.
D.1 Quantum Hypothesis Testing Protocol
The core idea of this protocol is to quantify the difficulty of distinguishing the target state $\rho$ from the reference state $\sigma_{R}$ through statistical experiments. In the framework of binary hypothesis testing ($H_{0}:\sigma_{R},H_{1}:\rho$), the type-I error is $\alpha=\mathrm{Tr}(T\sigma_{R})$, and the type-II error is $\beta=\mathrm{Tr}((I-T)\rho)$. The hypothesis testing relative entropy is defined as:
$D_{H}^{\eta}(\rho||\sigma_{R})=-\log\left(\min_{0\leq T\leq I,\mathrm{Tr}(T\sigma_{R})\leq\eta}\mathrm{Tr}((I-T)\rho)\right)$ (D.1)
The optimal test can be chosen from the family of threshold tests $T_{\gamma}=\{\rho-2^{\gamma}\sigma_{R}\geq 0\}$, by taking the smallest $\gamma$ that satisfies $\alpha(\gamma)\leq\eta$ (the quantum Neyman-Pearson lemma). However, any implementable test that satisfies $\alpha\leq\eta$ provides an operational lower bound for $D_{H}^{\eta}$.
Proposition D.1 (Confidence Lower Bound from Hypothesis Testing): By estimating $\beta$ through $N$ independent and repeated experiments and using the Clopper-Pearson (CP) method, one can construct a one-sided upper confidence endpoint $\beta_{U}(\delta)$ for $\beta$ at a confidence level of $1-\delta$. From this, a confidence lower bound for $D_{H}^{\eta}$ can be obtained:
$D_{H}^{\eta}(\rho||\sigma_{R})\geq-\log(\beta_{U}(\delta))$ (D.2)
In the "zero-failure" case, a sufficient condition for the minimum number of samples $N$ required to verify $D_{H}^{\eta}\geq L$ is:
$N\geq 2^{L}\cdot\ln(1/\delta)$ (D.3)
If $m$ tests or thresholds are scanned in the experiment, a multiple testing correction should be applied (e.g., Bonferroni: $\delta\mapsto\delta/m$, or Benjamini-Hochberg to control the FDR), and this should be noted in the report.
D.2 Projective Witness Protocol
The physical basis of this protocol is to "witness" the degree of information concentration in a state by measuring the occupation probability of a specific projection.
Lemma D.1 (Projective Witness Lower Bound): Let $\Pi$ be a projection of rank $r$ with $\mathrm{supp}(\Pi)\subseteq H_{R}$, and let $p=\mathrm{Tr}(\Pi\rho)$ be its occupation probability. Then the max-relative entropy $D_{\max}$ satisfies:
$D_{\max}(\rho||\sigma_{R})\geq\max\{0,\log(p\cdot d_{R}/r)\}$ (D.4)
The maximum value of $D_{\max}$ that can be certified by this method is $\log(d_{R}/r)$, corresponding to the case where $p=1$.
Proposition D.2 (Confidence Lower Bound from Projective Witnessing): By performing $N$ independent Bernoulli trials and using the CP method, one can construct a one-sided lower confidence endpoint $p_{L}(\delta)$ for $p$ at a confidence level of $1-\delta$. From this, a confidence lower bound for $D_{\max}$ can be obtained:
$D_{\max}(\rho||\sigma_{R})\geq\max\{0,\log(p_{L}(\delta)\cdot d_{R}/r)\}$ (D.5)
To verify a target $D_{\max}\geq L$ (equivalent to $p\geq p_{\mathrm{star}}=2^{L}\cdot r/d_{R}$), if there is an a priori estimate $p_{0}>p_{\mathrm{star}}$, then a sufficient condition for the required number of samples is given by the Hoeffding inequality:
$N_{\min}\geq\frac{1}{2(p_{0}-p_{\mathrm{star}})^{2}}\cdot\ln(1/\delta)$ (D.6)
If multiple sets of projections are scanned, a corresponding multiple testing correction must be performed.
D.3 Dephasing Entropy Bound Protocol
This protocol sets a conservative but robust bound on the complexity of a quantum state by measuring a "classicized" version of it. Consider the completely dephasing (pinching) channel $\Delta_{R}$ in a classical basis compatible with the reference set $R$, with $\Delta_{R}(\sigma_{R})=\sigma_{R}$. From the data processing inequality, we have:
$D(\rho||\sigma_{R})\geq D(\Delta_{R}(\rho)||\sigma_{R})=\log(d_{R})-H(P)$ (D.7)
where $P$ is the classical probability distribution corresponding to the dephased state, and $H(P)$ is its Shannon entropy.
Proposition D.3 (Confidence Upper Bound for Entropy): Let there be $M$ possible measurement outcomes. The plug-in entropy $H(P_{\mathrm{hat}})$ calculated from the empirical distribution $P_{\mathrm{hat}}$ obtained from $N$ samples is biased. To construct an upper confidence endpoint for the true entropy $H(P)$, at a confidence level of $1-\delta$, the L1 distance between $P_{\mathrm{hat}}$ and $P$ satisfies $P_{\mathrm{hat}}-P||_{1}\leq\varepsilon_{N}(\delta)$, where:
$\varepsilon_{N}(\delta)=\sqrt{\frac{2}{N}\cdot(M\ln 2+\ln(1/\delta))}$ (D.8)
Let $v=\varepsilon_{N}(\delta)/2$. According to the Fannes-Audenaert continuity bound for entropy, an upper confidence endpoint $H_{U}(\delta)$ for $H(P)$ can be obtained:
$H_{U}(\delta)=\min\{\log M,H(P_{\mathrm{hat}})+v\log(M-1)+h_{2}(v)\}$ (D.9)
where $h_{2}(v)$ is the binary Shannon entropy. This yields the final confidence lower bound:
$D(\rho||\sigma_{R})\geq\log(d_{R})-H_{U}(\delta)$ (D.10)
This appendix has provided complete implementation protocols with rigorous statistical control for the three operational principles proposed in the main text. These protocols ensure that the complexity lower bounds extracted from finite experimental or numerical data are not only operationally feasible but also statistically robust and reliable, thereby laying a solid foundation for the experimental verification and application of Theorem 3.1.
Appendix E Derivation and Physical Meaning of the RECT-$\eta$ Performance Equation
This appendix aims to provide a self-contained and complete derivation for the RECT-$\eta$ performance equation cited in Sec 5.2 of the main text, and to elucidate its nature as a “resource accounting identity”[31][60][61][32, 33].
E.1 Theoretical Foundations: Efficiency Factors and Equalization
The starting point of our theory is two independent, universal inequalities that describe the time cost of a quantum dynamical process; they constitute the axiomatic foundation of the RECT framework.
Bound from the Quantum Speed Limit (MT-type QSL-Complexity Lemma):
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}\geq\left(\frac{\pi\hbar}{2}\right)\cdot C_{\mathrm{opt}}$ (E.1)
Bound from the Principle of Locality (Locality-Entanglement Lemma):
$S_{E}\leq\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (E.2)
To transform these two bounds into exact expressions applicable to any process, we introduce two dynamical efficiency factors.
Definition E.1 (Efficiency Factors):
$\eta_{\mathrm{QSL}}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\sigma_{\mathrm{avail}}\cdot\Delta_{t}}$ (E.3)
$\eta_{\mathrm{LR}}=\frac{S_{E}}{(\gamma J/\hbar)\cdot\Delta_{t}}$ (E.4)
According to Eq. (E.1) and Eq. (E.2), the physical value ranges for these two efficiency factors are $\eta_{\mathrm{QSL}}\in(0,1]$ and $\eta_{\mathrm{LR}}\in[0,1]$, respectively. These two definitions precisely convert the aforementioned inequalities into equalities:
$\sigma_{\mathrm{avail}}\cdot\Delta_{t}=\frac{(\pi\hbar/2)\cdot C_{\mathrm{opt}}}{\eta_{\mathrm{QSL}}}$ (E.5)
$S_{E}=\eta_{\mathrm{LR}}\cdot\left(\frac{\gamma J}{\hbar}\right)\cdot\Delta_{t}$ (E.6)
E.2 Core Derivation
Eq. (E.5) and Eq. (E.6) provide two exact expressions for the same time interval $\Delta_{t}$ of the same physical process, which allows us to eliminate $\Delta_{t}$ through algebraic methods. Solving for $\Delta_{t}$ from both equations separately (assuming $S_{E}>0$):
$\Delta_{t}=\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}$ (E.7)
$\Delta_{t}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.8)
Setting the two expressions equal, we obtain:
$\frac{\pi\hbar C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{\hbar S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.9)
A profound physical phenomenon appears here: Planck’s constant $\hbar$ is completely canceled out on both sides of the equation. This is not a mathematical coincidence, but rather reflects an intrinsic duality wherein the time scales of both the QSL and the LR bound are calibrated by $\hbar$ in the same manner (linearly proportional). When comparing these two “cost ledgers” that have been calibrated by the same “ruler”, the unit of the ruler itself, $\hbar$, is naturally eliminated. After canceling out $\hbar$, we obtain a structural relationship that is independent of Planck’s constant:
$\frac{\pi C_{\mathrm{opt}}}{2\sigma_{\mathrm{avail}}\eta_{\mathrm{QSL}}}=\frac{S_{E}}{\eta_{\mathrm{LR}}\gamma J}$ (E.10)
Finally, by algebraic rearrangement to solve for the “resource-output product” $\sigma_{\mathrm{avail}}\cdot S_{E}$, we arrive at the RECT-$\eta$ equation:
$\sigma_{\mathrm{avail}}\cdot S_{E}=\left(\frac{\eta_{\mathrm{LR}}}{\eta_{\mathrm{QSL}}}\right)\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (E.11)
E.3 Physical Essence: The Identity and Testable Content
Proposition E.1 (The Identity Nature of RECT-$\eta$): Eq. (E.11), by its construction, is an algebraic identity.
Proof. Substitute the definitions of the efficiency factors, Eq. (E.3) and Eq. (E.4), into the right-hand side of Eq. (E.11):
$\mathrm{RHS}=\frac{S_{E}}{(\gamma J/\hbar)\Delta_{t}}\bigg/\frac{(\pi\hbar/2)C_{\mathrm{opt}}}{\sigma_{\mathrm{avail}}\Delta_{t}}\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$
After simplification, we obtain:
$\mathrm{RHS}$ $=\frac{S_{E}\cdot\sigma_{\mathrm{avail}}\cdot\Delta_{t}}{\bigl(\tfrac{\gamma J}{\hbar}\bigr)\Delta_{t}\cdot\bigl(\tfrac{\pi\hbar}{2}\bigr)C_{\mathrm{opt}}}\cdot\left(\frac{\pi\gamma J}{2}\right)\cdot C_{\mathrm{opt}}$ (E.12)
$=\frac{S_{E}\cdot\sigma_{\mathrm{avail}}}{\tfrac{\gamma J\pi\hbar^{2}}{2\hbar}}\cdot\left(\frac{\pi\gamma J}{2}\right)\qquad(\text{note: }\hbar^{2}/\hbar=\hbar)$ (E.13)
$=S_{E}\cdot\sigma_{\mathrm{avail}}.$ (E.14)
All constants, parameters, and $C_{\mathrm{opt}}$ are canceled out, ultimately yielding $\mathrm{RHS}=\sigma_{\mathrm{avail}}\cdot S_{E}$, which is equal to the left-hand side. Q.E.D. The correctness of this identity is not a falsifiable physical law. Its true, testable physical content is entirely contained within the two universal bounds that form its theoretical basis:
  1. 1.
    Physical Law 1 (QSL): $\eta_{\mathrm{QSL}}\leq 1$
  2. 2.
    Physical Law 2 (LR bound): $\eta_{\mathrm{LR}}\leq 1$
The value of the RECT-$\eta$ equation lies not in testing the equality itself, but in the fact that it provides a unified “resource accounting framework”. Within this framework, two seemingly unrelated physical laws are connected, allowing us to define a powerful set of diagnostic tools for dynamics (such as the efficiency ratio $\eta_{\mathrm{LR}}/\eta_{\mathrm{QSL}}$) and performance benchmarks (such as the performance frontier) to conduct in-depth analysis of real quantum processes.For the detailed derivation and discussion of this performance equation, please refer to the original paper [31].
Appendix F Generalized Settings and Framework Robustness
This appendix aims to explore the applicability of the Reference-Contingent Complexity (RCC) framework in more generalized settings and to address profound questions regarding the objectivity of its core constructs. Section F.1 generalizes the theory to reference sets defined by non-Abelian symmetries. Section F.2 provides a rigorous scheme for a conservative lower bound for handling systems that require an energy or mode cutoff (such as quantum field theories). Finally, Section F.3 provides physical, objectivization criteria for the selection of the reference set and proves the robustness of the framework with respect to this choice[62, 63].
F.1 Non-Abelian Reference Sets
The theoretical construction in the main text is primarily based on Abelian-type reference sets defined by mutually commuting projectors. However, the framework can be naturally generalized to systems defined by non-Abelian group symmetries.
Definition F.1 (Non-Abelian Reference Set): Let the Hilbert space $H$ of a system, under the action of a compact group $G$, be decomposed into a direct sum of irreducible representations (irreps) according to the Peter-Weyl theorem:
$H\cong\bigoplus_{\lambda\in\hat{G}}(V_{\lambda}\otimes M_{\lambda})$
where $V_{\lambda}$ is the $d_{\lambda}$-dimensional irrep space, and $M_{\lambda}$ is the $m_{\lambda}$-dimensional multiplicity space. We select a subset of irreps $\Lambda\subseteq\hat{G}$ and, in each corresponding multiplicity space $M_{\lambda}$, select a projector $Q_{\lambda}$ of rank $r_{\lambda}\leq m_{\lambda}$. The total projector $\Pi_{R}$ of the non-Abelian reference set and its reference state $\sigma_{R}$ are defined as:
$\Pi_{R}=\bigoplus_{\lambda\in\Lambda}(I_{V_{\lambda}}\otimes Q_{\lambda})\\ d_{R}=\mathrm{Tr}(\Pi_{R})=\sum_{\lambda\in\Lambda}d_{\lambda}\cdot r_{\lambda}\\ \sigma_{R}=\Pi_{R}/d_{R}$ (F.1)
Lemma F.1 (Relative Entropy in the Non-Abelian Setting): If $\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\Pi_{R})$, then the expressions for $D(\rho||\sigma_{R})$ and $D_{\max}(\rho||\sigma_{R})$ are identical to the Abelian case:
$D(\rho||\sigma_{R})=\log(d_{R})-S(\rho)$ (F.2)
$D_{\max}(\rho||\sigma_{R})=\log(d_{R})-H_{\min}(\rho)$ (F.3)
Corollary F.1 (Entropy Identity and Monotonicity): In the non-Abelian setting, the entropy identity $D_{\max}-D=S-H_{\min}$ still holds. Concurrently, for any cluster-preserving channel $E$ that leaves the reference state invariant (i.e., a CPTP map satisfying $E(\sigma_{R})=\sigma_{R}$), the monotonicity property $C_{R}(E(\rho))\leq C_{R}(\rho)$ also holds.
F.2 Conservative Lower Bound under Energy or Mode Cutoff
When dealing with systems that have an infinite-dimensional Hilbert space, any numerical or experimental analysis must introduce an energy or mode cutoff $\Lambda$. We prove here that the RCC framework can provide a rigorous, conservative lower bound on complexity in this setting. Let $\Pi_{[0,\Lambda]}$ be the projection operator that preserves modes with energy below $\Lambda$. We define a truncated reference set $\Pi_{R}^{(\Lambda)}=\Pi_{[0,\Lambda]}\cdot\Pi_{R}$, with its corresponding reference state being $\sigma_{R}^{(\Lambda)}$. To avoid support issues, we employ a smoothed reference state $\tilde{\sigma}_{R}^{(\Lambda)}=(1-\delta)\sigma_{R}^{(\Lambda)}\oplus\delta\tau_{\perp}$, where $\tau_{\perp}$ is the equally-mixed state on the orthogonal complement of $\Pi_{R}^{(\Lambda)}$. For a physical state $\rho_{\mathrm{phys}}$, its retention probability within the truncated subspace is $q=\mathrm{Tr}(\Pi_{R}^{(\Lambda)}\cdot\rho_{\mathrm{phys}})$, and the normalized truncated state is $\rho^{(\Lambda)}$.
Lemma F.2 (Truncation Error Bound): For a "measure-and-prepare" channel that measures "whether the state is within $\Pi_{R}^{(\Lambda)}$", from the data processing inequality for relative entropy and its decomposition, we have:
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq D_{\mathrm{Bern}}(q||1-\delta)+q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (F.4)
where $D_{\mathrm{Bern}}$ is the classical Bernoulli relative entropy (base 2).
Corollary F.2 (Multiplicative Reduction Lower Bound): By choosing the smoothing parameter $\delta=1-q$, the $D_{\mathrm{Bern}}$ term becomes zero, and we obtain a concise multiplicative reduction bound:
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (F.5)
This expression clearly shows that when an energy cutoff or state leakage is present, the leading term of the measurable complexity lower bound is linearly reduced by the retention probability $q$ of the state within the target subspace.
F.3 Acceptability and Objectivity of the Reference Set
The effectiveness of the RCC framework depends on the selection of a physically meaningful reference set $R$. To eliminate potential subjective arbitrariness in its selection, we propose the following criteria and robustness characterization. A physically acceptable reference set should satisfy the following criteria:
  1. 1.
    Structural Invariance: All dynamical processes in the system that are considered "free" or "zero-cost" should preserve the structure of the reference set.
  2. 2.
    Measurability: There must exist physical measurement schemes compatible with $R$ that can effectively estimate any of the core information quantities in Appendix D.
  3. 3.
    Principle of Minimal Sufficiency: Among the candidates that satisfy the first two criteria, one should preferentially choose the "coarsest" structure (i.e., the one with the largest $d_{R}$), so as to avoid overestimating the complexity.
Proposition F.1 (Misspecification Robustness Bound): Suppose a state $\rho$ is supported by two different reference sets, $R$ and $R^{\prime}$. The difference between their corresponding complexities, $C_{R}(\rho)$ and $C_{R^{\prime}}(\rho)$, is explicitly constrained by the parameters of the reference sets themselves:
$|C_{R}(\rho)-C_{R^{\prime}}(\rho)|\leq\left|\frac{\log(d_{R})}{\log(\Gamma_{R})}-\frac{\log(d_{R^{\prime}})}{\log(\Gamma_{R^{\prime}})}\right|+\left|\frac{1}{\log(\Gamma_{R})}-\frac{1}{\log(\Gamma_{R^{\prime}})}\right|\cdot S(\rho)$ (F.6)
(This bound only uses $(d_{R},\Gamma_{R})$ and $S(\rho)$; for the same $\rho$, this deviation is computable and controllable.)
In particular, if the hardware capability is unchanged ($\Gamma_{R}=\Gamma_{R^{\prime}}$) and $R^{\prime}$ is a refinement of $R$ ($d_{R^{\prime}}<d_{R}$), then there is an exact difference relation:
$0\leq C_{R}(\rho)-C_{R^{\prime}}(\rho)=\frac{\log(d_{R})-\log(d_{R^{\prime}})}{\log(\Gamma_{R})}$ (F.7)
This proposition transforms the effect of "subjective selection of a reference set" into a quantitatively computable and controllable deviation, thereby mathematically proving the robustness of the framework.
Appendix G Near-Tightness of Reference-Contingent Complexity
This appendix aims to provide constructive, non-trivial, positive evidence for Conjecture 6.1 (the near-tightness of $RCC$), which was proposed in Section 6.2 of the main text. The conjecture posits that Reference-Contingent Complexity $C_{R}(\rho)$ is not only a lower bound on the universal circuit complexity $C_{\mathrm{opt}}(\rho)$, but that for a class of well-structured families of states, it can also be bounded from above by $C_{R}(\rho)$. We will systematically corroborate this conjecture by constructing complexity upper bounds for three core families of states that are widely present in quantum information and physics: efficiently samplable classical diagonal states, low-rank states constructible via hierarchical projections, and efficiently encodable logical pure states with sparse generators. These constructive proofs collectively reveal that for these physically important states, the scaling behavior of $C_{\mathrm{opt}}(\rho)$ is indeed closely related to that of $C_{R}(\rho)$, indicating that $C_{R}(\rho)$ is not merely a theoretical baseline, but a physical quantity with strong predictive power for real computational resources.
G.1 Technical Conventions and Parameterization
To ensure the rigor of the arguments, we first unify the technical conventions of this appendix. All logarithms $\log$ are taken to base $2$. The error metric used is the trace distance, $(1/2)\cdot||.||_{1}$. All core physical quantities in this paper, such as Reference-Contingent Complexity $C_{R}(\rho)$, the reference state $\sigma_{R}$, the subspace dimension $d_{R}$, and the information bandwidth parameter $\Gamma_{R}$, follow the definitions in Chapter 2 of the main text. According to the Stinespring dilation theorem [64], any completely positive and trace-preserving (CPTP) map can be realized by a unitary evolution on a larger Hilbert space, followed by tracing out an ancillary system. This fundamental principle ensures that all seemingly non-unitary constructions in this appendix (such as measurements, random mixing) can be compiled into an equivalent, post-selection-free quantum circuit, with the constant factors in their costs being uniformly absorbed into the final upper bound constants. A crucial clarification concerns the parameter dependence of the upper bound constants. To avoid circular logic, we must precisely define the scope of independence for these constants.
Definition G.1 (Parameterized Upper Bound Constants for Complexity). In Conjecture 6.1 and all upper bound constructions in this appendix of the form $C_{\mathrm{opt}}\leq a\cdot C_{R}+b\cdot(\dots)$, the constants $a$ and $b$ are assumed to be $O(1)$ quantities that are independent of the system size $d_{R}$ and the specific target state $\rho$. However, they are allowed to depend on the information bandwidth parameter $\Gamma_{R}$, which is uniquely determined by the hardware architecture (i.e., the universal gate set $g$ and the addressability $|S_{R}|$). Therefore, the rigorous form of the conjecture should be written as:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.1)
This parameterization both ensures the validity of the core assertion (that the scaling behavior is independent of $d_{R}$) and aligns it with the intrinsic computational capabilities of different physical platforms.
G.2 Case One: Efficiently Samplable Classical Diagonal States
We first consider a class of states that are of paramount importance in statistical physics and machine learning: their density matrices are diagonal in a computational basis $\{|x\rangle\}$ that is compatible with the reference set $R$, and thus are physically equivalent to a classical probability distribution.
Definition G.2 (Family of Classical Diagonal States). A classical diagonal state $\rho$ is a quantum state of the form $\rho=\sum_{x}p_{x}|x\rangle\langle x|$, where $p=\{p_{x}\}$ is a probability distribution over $d_{R}=2^{n}$ basis states, and $H(p)=-\sum_{x}p_{x}\log(p_{x})$ is its Shannon entropy.
For an arbitrary probability distribution $p$, the cost of preparing the corresponding quantum state $\rho$ can be extremely high. However, if the distribution satisfies certain conditions of efficient samplability, its complexity is greatly simplified.
Assumption G.1 (Efficient Amplitude Preparation). A probability distribution $p$ is considered efficiently preparable in amplitude if there exists a unitary circuit $U_{\mathrm{amp}}$, acting on $n$ ancillary qubits, that can prepare its amplitude-encoded state $|\psi_{p}\rangle=\sum_{x}\sqrt{p_{x}}|x\rangle$ within a precision $\varepsilon^{\prime}$ in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{amp}})\leq c_{1}\cdot n+c_{2}\cdot\log(1/\varepsilon^{\prime})$. This assumption covers an important class of probability distributions in physics and algorithms, for which sufficient conditions include, but are not limited to: (i) the distribution can be described by an autoregressive model, and its conditional probabilities $p(x_{i}|x_{<i})$ can be efficiently computed classically; (ii) the distribution comes from a probabilistic graphical model with low tree-width [65, 66, 67, 68].
Proposition G.1. For a classical diagonal state $\rho$ that satisfies Assumption G.1, its optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ satisfies the following upper bound:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.2)
Proof (Constructive). We can prepare $\rho$ via a three-step, post-selection-free unitary circuit. Consider a composite system consisting of a main system $S$ ($n$-qubit) and an ancillary system AA ($n$-qubit), with the initial state being $|0...0\rangle_{S}\otimes|0...0\rangle_{A}$.
  1. 1.
    Amplitude Preparation: On the ancillary system $A$, execute the unitary operation $U_{\mathrm{amp}}$ to transform the system into $|0...0\rangle_{S}\otimes|\psi_{p}\rangle_{A}$. According to Assumption G.1, the cost of this step is $O(n)=O(\log d_{R})$.
  2. 2.
    Controlled Copying: Sequentially, for i=1ni=1...n, perform a CNOT operation with the $i$-th qubit of $A$ as the control and the $i$-th qubit of $S$ as the target. This copies the classical information from $|\psi_{p}\rangle_{A}$ into $S$, resulting in the entangled state $\sum_{x}\sqrt{p_{x}}|x\rangle_{S}\otimes|x\rangle_{A}$. This step consists of $n$ CNOT gates, and its serial cost is $O(n)=O(\log d_{R})$.
  3. 3.
    Decoupling: (Conceptually) Trace out the ancillary system $A$. Since $\mathrm{Tr}_{A}[(\sum_{x}\sqrt{p_{x}}|x\rangle_{S}\otimes|x\rangle_{A})(\sum_{y}\sqrt{p_{y}}\langle y|_{S}\otimes\langle y|_{A})]=\sum_{x}p_{x}|x\rangle_{S}\langle x|_{S}=\rho$, we have successfully prepared the target state $\rho$ on the main system $S$.
The entire process is a deterministic CPTP map, and the total upper bound on the number of atomic instruction steps is $O(\log d_{R})+O(\log(1/\varepsilon))$, which proves Eq. (G.2). This result is in high agreement with Conjecture 6.1. $C_{R}(\rho)$ is determined by $(\log d_{R}-H(p))$, while the upper bound on $C_{\mathrm{opt}}(\rho)$ is dominated by $\log d_{R}$. The scaling behavior of the two is consistent when $H(p)$ is much smaller than $\log d_{R}$, and when $H(p)$ approaches $\log d_{R}$, the term $b\cdot\log d_{R}$ provides the correct "fallback" upper bound.
G.3 Case Two: Low-Rank, High-Occupation States and Hierarchical Projection Construction
The second important family of states are those whose information is highly concentrated in a small-dimensional subspace of the Hilbert space.
Definition G.3 (Family of Low-Rank, High-Occupation States). A low-rank, high-occupation state $\rho$ is a quantum state of the form $\rho=p\cdot(\Pi/r)+(1-p)\cdot\tau$, where $\Pi$ is a projection operator of rank $r\ll d_{R}$, $p$ is a high occupation probability close to $1$, and $\tau$ is a background state supported within $H_{R}$.
Assumption G.2 (Hierarchical Projectability). A projector $\Pi$ is considered hierarchically constructible if it can be defined by the product $\Pi=\prod_{j}P_{j}$ of $m=O(\log(d_{R}/r))$ mutually commuting projection operators $\{P_{j}\}$. This is a common structure in stabilizer codes and ground states of commuting local Hamiltonians.
Proposition G.2. For a low-rank, high-occupation state $\rho$ that satisfies Assumption G.2, its optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ satisfies the following upper bound:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.3)
Proof (Constructive). Our core is to construct a CPTP channel $T_{\Pi}$ that deterministically maps the maximally mixed reference state $\sigma_{R}$ to the mixed state on the target subspace, $\Pi/r$.
Lemma G.2.1. There exists a CPTP channel $T_{\Pi}$ such that $T_{\Pi}(\sigma_{R})=\Pi/r$, and its circuit implementation cost is $O(m)=O(\log(d_{R}/r))$.
Proof of Lemma. We can construct $T_{\Pi}$ by concatenating $m$ "measure-reprepare-depolarize" quantum instruments $\Phi^{(j)}$. For each $P_{j}$, define $\Phi^{(j)}(X)=P_{j}\cdot X\cdot P_{j}+\mathrm{Tr}[(I-P_{j})\cdot X]\cdot(P_{j}/\mathrm{Tr}(P_{j}))$. This is a post-selection-free CPTP map, which physically corresponds to measuring the eigenvalue of $P_{j}$: if the outcome is +1+1, the state is kept; if the outcome is 1-1, that part of the state is reinjected into the +1+1 subspace of $P_{j}$. Since all $P_{j}$ commute, these $m$ channels can be applied sequentially, with each step compressing the state into a smaller subspace, ultimately mapping $\sigma_{R}$ exactly to $\Pi/r$. The implementation cost of each step $\Phi^{(j)}$ depends on the structure of $P_{j}$. If $P_{j}$ acts on a local subsystem, this operation can be realized via a local $2$-design with $O(1)$ cost [69, 70]. If $P_{j}$ is non-local, its implementation cost is absorbed as a constant factor into the total cost of $O(\log(d_{R}/r))$. Using this lemma, we can prepare $\rho$ by introducing a control qubit: with probability $p$, apply the channel $T_{\Pi}$ to $\sigma_{R}$, and with probability 1p1-p, prepare the background state $\tau$. The total cost is dominated by $O(\log(d_{R}/r))$ and the cost of preparing $\tau$, with an upper bound of $O(\log d_{R})$. On the other hand, from the convexity of relative entropy, the leading term of $C_{R}(\rho)$ is determined by $p\cdot\log(d_{R}/r)$. Therefore, the upper bound on $C_{\mathrm{opt}}(\rho)$ is consistent in scaling with the leading term of $C_{R}(\rho)$, thus verifying Eq. (G.3).
G.4 Case Three: Sparse Graph States and Efficiently Encodable Logical Pure States
The final class, and the one that most strongly demonstrates the near-tightness of $RCC$, comprises the core objects of quantum computation and error correction theory—logical pure states.
Definition G.4 (Family of Efficiently Encodable Logical Pure States). A logical pure state $\rho=|\psi_{L}\rangle\langle\psi_{L}|$ is a pure state located within a $d_{R}=2^{k}$-dimensional code subspace $H_{R}$. It is considered efficiently encodable if there exists an encoding unitary circuit $U_{\mathrm{enc}}$ that can map a simple product state $|s_{0}\rangle$ to a logical ground state $|0\rangle_{L}$ within the code space, in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{enc}})\leq c\cdot k$. Logical operations can then be performed on this state. Typical representatives of this class of states are: (i) graph states with bounded vertex degree ($|E|=O(k)$); (ii) logical states of stabilizer codes that are geometrically sparse or satisfy LDPC conditions [15, 71].
Proposition G.3. For a logical pure state $\rho$ belonging to an efficiently encodable family, its optimal circuit complexity $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ is directly proportional to its Reference-Contingent Complexity $C_{R}(\rho)$:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot\log(1/\varepsilon)$ (G.4)
Proof (Direct Construction). For this class of states, the upper bound on $C_{\mathrm{opt}}(\rho)$ is directly given by the most efficient encoding circuit, i.e., $C_{\mathrm{opt}}(\rho)\leq O(k)=O(\log d_{R})$. On the other hand, since $\rho$ is a pure state, its von Neumann entropy is $S(\rho)=0$. Therefore, its Reference-Contingent Complexity is:
$C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}}=\frac{k}{\log\Gamma_{R}}$ (G.5)
Comparing the upper bound on $C_{\mathrm{opt}}(\rho)$, which is $O(k)$, with $C_{R}(\rho)=k/\log\Gamma_{R}$, we find that they differ only by a factor of $a(\Gamma_{R})=O(\log\Gamma_{R})$, and their scaling behaviors are identical. This directly proves Eq. (G.4), providing the strongest corroborating evidence for Conjecture 6.1. Conclusion: Through a constructive analysis of three families of states that are physically and computationally distinct but are all "well-structured," we have consistently found that the upper bound on their optimal preparation complexity, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, follows the form $C_{\mathrm{opt}}\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$. This series of non-trivial pieces of evidence strongly supports the validity of Conjecture 6.1 proposed in the main text. This profoundly suggests that Reference-Contingent Complexity $C_{R}(\rho)$ is not just an abstract, possibly loose, theoretical lower bound, but is, for a broad class of physically relevant states, a "near-tight" physical benchmark that can effectively estimate the true computational complexity.
Appendix H Notation and Core Terminology
This appendix provides a centralized quick reference list for the core symbols, key terms, and their abbreviations as defined and used in the main text. All entries are categorized according to their logical function in the paper to facilitate reader reference.
H.1 Core Complexity Measures
$C_{R}(\rho)$
Reference-Contingent Complexity (RCC). Defined as $D(\rho||\sigma_{R})/\log\Gamma_{R}$, which is the quantum relative entropy of the state $\rho$ with respect to its structured vacuum $\sigma_{R}$, normalized by the information bandwidth.
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$
Optimal Circuit Complexity. Under the given computational model, the minimum number of serial atomic instruction steps required to prepare the quantum state $\rho$ to a precision $\varepsilon$ (in trace distance).
$K_{\varepsilon}(\rho|R)$
Smoothed Conditional Kolmogorov Complexity. On a universal Turing machine conditioned on a description of the reference set $R$, the length (in bits) of the shortest program that can output the quantum state $\rho$ within a precision $\varepsilon$.
H.2 Quantum States and Operators
$\rho$
Target Quantum State. Typically a density matrix, it is the object of the complexity analysis.
$\sigma_{R}$
Reference State. The maximally mixed state on the subspace $H_{R}$ defined by the reference set $R$, physically playing the role of a ’structured vacuum’.
$\Pi_{R}$
Total Projector. The projection operator formed by the product $\prod_{i}P_{i}$ of all constraint projectors $\{P_{i}\}$ in the reference set $R$, whose image space is the structured subspace $H_{R}$.
$P_{i}$
Set of Constraint Projectors. A set of mutually commuting projection operators that together constitute the reference set $R$, formalizing the structured priors of the system.
$H_{R}$
Structured Subspace. The Hilbert subspace defined by the total projector $\Pi_{R}$, $H_{R}=\Pi_{R}\cdot H$.
H.3 Information Theory and Entropic Quantities
$S(\rho)$
von Neumann Entropy. Defined as $\mathrm{Tr}(\rho\log\rho)$, it quantifies the degree of mixedness of the quantum state $\rho$.
$D(\rho||\sigma)$
Quantum Relative Entropy. Defined as $\mathrm{Tr}[\rho(\log\rho-\log\sigma)]$, it quantifies the distinguishability between two quantum states $\rho$ and $\sigma$.
$D_{\max}(\rho||\sigma)$
Max-Relative Entropy. Defined as $\inf\{\lambda:\rho\leq 2^{\lambda}\cdot\sigma\}$, it is a core measure in one-shot information theory.
$H_{\min}(\rho)$
[…]
$\Delta_{\mathrm{spec}}(\rho)$
Spectral Skewness Term. Defined as $S(\rho)-H_{\min}(\rho)$, it quantifies the non-uniformity of the spectral distribution of the quantum state $\rho$.
$D_{H}^{\eta}(\rho||\sigma)$
Hypothesis Testing Relative Entropy. In the task of distinguishing $\rho$ from $\sigma$, it represents the minimum achievable exponential decay rate of the type-II error probability, given that the type-I error probability does not exceed $\eta$.
H.4 Model Parameters
$d_{R}$
Subspace Dimension. The dimension of the structured subspace $H_{R}$, $d_{R}=\mathrm{Tr}(\Pi_{R})$.
$\Gamma_{R}$
Information Bandwidth Parameter. Defined as $g\cdot|S_{R}|$, i.e., the size of the alphabet of atomic instructions. $\log\Gamma_{R}$ is the information cost of a single atomic instruction.
$g$
Gate Set Cardinality. The total number of distinct unitary gates in the universal gate set $G$.
$|S_{R}|$
Number of Addressable Units. The total number of target locations or units on which a single atomic instruction can act.
$\varepsilon$
Precision. The maximum acceptable trace distance between the prepared quantum state and the ideal target state.
Atomic Instruction Alphabet. The set composed of all possible atomic instructions $(u,s)$, whose size is $\Gamma_{R}$.
[…]
H.5 Core Terminology
Atomic Instruction
The fundamental operational unit of the computational model, consisting of an ordered pair $(u,s)$ composed of a unitary gate $u$ and a target of action $s$.
Reference Cluster
A set composed of mutually commuting constraint projectors, $R=\{P_{i}\}$, used to mathematically formalize the intrinsic structure and priors of a system.
Structural Fairness
An ideal property that a complexity measure should possess, meaning that the measure only prices the non-trivial information that transcends the system’s intrinsic ’free’ structure.
Appendix I Complete Derivation of the “Work–Complexity” and “Time–Complexity” Relations
This appendix aims to provide a complete and rigorous mathematical derivation and boundary condition analysis for the “Work–Complexity” identity (Proposition 5.3) and the “Process Time Lower Bound” (Theorem 5.2) proposed in Sec. 5.4 of the main text. To ensure the rigor of the discussion and consistency of the framework, we first clarify the following conventions: in this appendix, unless explicitly specified as $\log_{2}$, all instances of $\log$ refer to the natural logarithm $\ln$, and entropy is measured in nats, so as to maintain consistency with the thermodynamic framework. When citing content from the main text, a standard unit conversion between $\mathrm{bits}\leftrightarrow\mathrm{nats}$ will be assumed[35][72].
I.1 Core Setting
To proceed with the subsequent derivations, we first establish the core mathematical and physical setting, which specifically includes the following four aspects:
  1. 1.
    System and Reference Set: A finite-dimensional Hilbert space $H$. The reference set $R$ is defined by a set of mutually commuting projectors $\{P_{i}\}$, with the total projector being $\Pi_{R}=\prod_{i}P_{i}\neq 0$. The subspace is $H_{R}=\Pi_{R}H$ (with dimension $d_{R}=\mathrm{Tr}(\Pi_{R})$), and the reference state is $\sigma_{R}=\Pi_{R}/d_{R}$.
  2. 2.
    Windowed Coarse-Graining: We adopt the conditional expectation maps $\{E_{\Xi}\}$ ($\Xi\geq 0$) from CWT [35], which correspond to a nested algebra of observables $\{A(\Xi)\}$ satisfying
    $E_{\Xi}(\sigma_{R})=\sigma_{R},\quad E_{\Xi}\circ E_{\Xi}=E_{\Xi},\quad\Xi_{1}\leq\Xi_{2}\Rightarrow A(\Xi_{1})\subseteq A(\Xi_{2})$ (I.1.1)
    The windowed entropy is defined as $S_{\Xi}(\rho)=S(E_{\Xi}(\rho))$, where $S$ is the von Neumann entropy.
  3. 3.
    CWT Thermodynamic Accounting: The energy is $E=E(S_{\Xi},\Xi,\dots)$; its differential is $\mathrm{d}E=T_{\Xi}\mathrm{d}S_{\Xi}+\Pi_{\Xi}\mathrm{d}\Xi+\dots$, where $T_{\Xi}=(\partial E/\partial S_{\Xi})_{\Xi}$ and $\Pi_{\Xi}=(\partial E/\partial\Xi)_{S_{\Xi}}$. Under a fixed-energy constraint, $\Pi_{\Xi}=-T_{\Xi}\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$.
  4. 4.
    RCC Definitions: The information bandwidth $\Gamma_{R}$ is determined by $R$ and the hardware addressability (see Sec. 2.1). The standard and windowed RCC are:
    $C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}},\quad C_{R,\Xi}(\rho)=\frac{\log d_{R}-S_{\Xi}(\rho)}{\log\Gamma_{R}}$ (I.1.2)
    It is assumed by default that $\Gamma_{R}$ is independent of $\Xi$; for a generalization, see Sec. I.3.2. Relevant functions are assumed to have sufficient smoothness.
I.2 Mathematical Properties of Windowed RCC
Based on the setting above, the windowed RCC ($C_{R,\Xi}$), as a core measure, exhibits the following key mathematical properties:
Proposition I.1 (Monotonicity and Limits). Under the setting (I.1.1), for any state $\rho$:
  1. (i)
    If $\Xi_{1}\leq\Xi_{2}$, then $S_{\Xi_{1}}(\rho)\geq S_{\Xi_{2}}(\rho)$, and therefore $C_{R,\Xi}$ is monotonically non-decreasing with $\Xi$;
  2. (ii)
    $\lim_{\Xi\to\infty}S_{\Xi}(\rho)=S(\rho)$, hence $\lim_{\Xi\to\infty}C_{R,\Xi}(\rho)=C_{R}(\rho)$.
Proof. From $S_{\Xi}(\rho)=\log d_{R}-D(E_{\Xi}(\rho)||\sigma_{R})$, $E_{\Xi_{1}}=E_{\Xi_{1}}\circ E_{\Xi_{2}}$, and $E_{\Xi}(\sigma_{R})=\sigma_{R}$, applying the data processing inequality to the relative entropy yields $D(E_{\Xi_{1}}(\rho)||\sigma_{R})\leq D(E_{\Xi_{2}}(\rho)||\sigma_{R})$; hence the conclusion holds.
I.3 The "Work–Complexity" Relation: From Identity to Boundary Cases
This section aims to establish a precise mathematical relationship between the generalized force (the differential form of work) and the change in complexity.
I.3.1 Identity and Work Accounting in the Standard Case
Theorem I.1 (Work–Complexity Identity).
$\Pi_{\Xi}=T_{\Xi}\cdot\log(\Gamma_{R})\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}$ (I.3.1)
Proof. From (I.1.2), we have $(\partial C_{R,\Xi}/\partial\Xi)_{E}=-(1/\log\Gamma_{R})\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$. Using $\Pi_{\Xi}=-T_{\Xi}\cdot(\partial S_{\Xi}/\partial\Xi)_{E}$ then yields the result.
Corollary I.1 (Rigorous Work Accounting and Weighted Average Temperature).
$W_{\mathrm{info}}=\int_{\Xi_{i}}^{\Xi_{f}}\Pi_{\Xi}\mathrm{d}\Xi=\log\Gamma_{R}\cdot\int_{C_{R,\Xi_{i}}}^{C_{R,\Xi_{f}}}T_{\Xi}\mathrm{d}C_{R,\Xi}=\log\Gamma_{R}\cdot\langle T_{\Xi}\rangle_{C}\cdot\Delta C_{R,\Xi}$ (I.3.2)
where $\langle T_{\Xi}\rangle_{C}=\left(\int T_{\Xi}\mathrm{d}C_{R,\Xi}\right)/\Delta C_{R,\Xi}$.
I.3.2 Generalization: Variable Bandwidth $\Gamma_{R}(\Xi)$
When $C_{R,\Xi}=(\log d_{R}-S_{\Xi})/\log\Gamma_{R}(\Xi)$, the chain rule gives
$\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}=-\frac{1}{\log\Gamma_{R}}\cdot\left(\frac{\partial S_{\Xi}}{\partial\Xi}\right)_{E}-\frac{C_{R,\Xi}}{\log\Gamma_{R}}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)$ (I.3.3)
Substituting this into the definition of $\Pi_{\Xi}$ yields the generalized identity
$\Pi_{\Xi}=T_{\Xi}\cdot\left[\log\Gamma_{R}\cdot\left(\frac{\partial C_{R,\Xi}}{\partial\Xi}\right)_{E}+C_{R,\Xi}\cdot\left(\frac{\partial\log\Gamma_{R}}{\partial\Xi}\right)\right]$ (I.3.4)
I.4 The "Time–Complexity" Lower Bound: From Strict to Complete Form
This section derives the lower bound on the time required to complete a computational task, a bound determined by the change in complexity.
I.4.1 Strict Lower Bound and Physical Form
CWT provides an upper bound on the envelope rate, $\dot{C}^{*}(t)\leq(2/\hbar)\Pi_{\Xi}(t)$ (almost everywhere). Integrating over time yields
$C^{*}(t_{f})-C^{*}(0)\leq\frac{2}{\hbar}\cdot t_{f}\cdot\langle\Pi_{\Xi}\rangle_{t}$ (I.4.1)
Combining this with the main theorem, $C^{*}(t_{f})\geq C_{R}(\rho_{f})$, gives
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)]$ (I.4.2)
Under the conditions of a compact initial state or a non-retreating envelope, this simplifies to
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C_{R}(\rho_{i})]$ (I.4.3)
For an isothermal process and taking $\Pi_{\Xi}\lesssim\pi k_{B}T_{\Xi}$, we get
$t_{f}\gtrsim\frac{\hbar}{2\pi k_{B}\bar{T}_{\Xi}}\cdot\Delta C_{R}$ (I.4.4)
I.4.2 Lower Bound with Full Correction Terms
Substituting the full form of the main theorem, let
$C_{R}^{\mathrm{full}}(\rho)=C_{R}(\rho)+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}\\ C_{\log}(t,\varepsilon)=\frac{c_{1}}{\log\Gamma_{R}}\cdot\left[\log(\max\{1,C^{*}(t)\})+\log(1/\varepsilon)\right]$ (I.4.5)
This gives
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}\rangle_{t}}\cdot[C_{R}^{\mathrm{full}}(\rho_{f})-C_{R}^{\mathrm{full}}(\rho_{i})-\Delta C_{\log}^{\mathrm{upper}}]$ (I.4.6)
where $\Delta C_{\log}^{\mathrm{upper}}=C_{\log}(t_{f},\varepsilon)-C_{\log}(0,\varepsilon)$ is the difference in the logarithmic correction terms. If it is necessary to eliminate the bootstrap nature of this bound to obtain a closed-form lower bound, one can refer to the Lambert-W technique in Appendix C.2.
I.4.3 Boundary Case: $\Pi_{\Xi}$ Changing Sign
If $\Pi_{\Xi}<0$ during some part of the process, considering $\Pi_{\Xi}^{+}(t)=\max\{\Pi_{\Xi}(t),0\}$ yields a sign-robust lower bound
$t_{f}\geq\frac{\hbar}{2\langle\Pi_{\Xi}^{+}\rangle_{t}}\cdot[C_{R}(\rho_{f})-C^{*}(0)],\quad\langle\Pi_{\Xi}^{+}\rangle_{t}=\frac{1}{t_{f}}\int_{0}^{t_{f}}\Pi_{\Xi}^{+}(t)\mathrm{d}t$ (I.4.7)
I.5 Robustness Analysis of the Framework
Finally, we analyze the robustness of this theoretical framework in the face of two typical practical challenges: leakage of the physical state and statistical uncertainty in the measures.
I.5.1 Handling of Support Set Leakage
If the physical state $\rho$ has a small leakage $p_{\mathrm{leak}}$ with respect to the target subspace $H_{R}$, the projected-and-renormalized state $\tilde{\rho}$ has a trace distance
[…]
By the Fannes–Audenaert continuity bound,
$|S_{\Xi}(\tilde{\rho})-S_{\Xi}(\rho)|\leq\delta\log(d_{R}-1)+h_{2}(\delta)$ (I.5.2)
which implies
$|C_{R,\Xi}(\tilde{\rho})-C_{R,\Xi}(\rho)|\leq\frac{(\delta\log(d_{R}-1)+h_{2}(\delta))}{\log\Gamma_{R}}$ (I.5.3)
When $p_{\mathrm{leak}}$ is sufficiently small, $\delta=O(\sqrt{p_{\mathrm{leak}}})$, and the error is controllable.
I.5.2 Statistical Ledger for Operational Estimation
Following the protocol in Chapter 4, one can obtain a lower bound $L_{H}$ for $D_{H}^{\eta}(E_{\Xi}(\rho)||\sigma_{R})$ at a confidence level of $1-\delta$, which in turn gives a statistical lower bound for the main term of $C_{R,\Xi}$, $C_{R,\Xi}^{\mathrm{(main)}}\geq L_{H}/\log\Gamma_{R}$. Substituting this into (I.3.2) and (I.4.2) allows for the estimation of confidence intervals for $W_{\mathrm{info}}$ and $t_{f}$.